### v2

# Sard properties for polynomial maps in infinite dimension

Abstract.
Sard’s theorem asserts that the set of critical values of a smooth map from one Euclidean space to another one has measure zero. A version of this result for infinite-dimensional Banach manifolds was proven by Smale for maps with Fredholm differential. It is well–known, however, that when the domain is infinite dimensional and the range is finite dimensional, the result is not true – even under the assumption that the map is “polynomial” – and a general theory is still lacking. Addressing this issue, in this paper, we provide sharp quantitative criteria for the validity of Sard’s theorem in this setting. Our motivation comes from sub–Riemannian geometry and, as an application of our results, we prove the sub–Riemannian Sard conjecture for the restriction of the Endpoint map of Carnot groups to the set of piece–wise real–analytic controls with large enough radius of convergence, and the strong Sard conjecture for the restriction to the set of piece–wise entire controls.
1. Introduction
1.1. The Sard property for polynomial maps on infinite-dimensional spaces
The classical Sard theorem [Mor39, Sar42] states that the set of critical values of a smooth map $f:N\to M$, where $N$ and $M$ are finite dimensional manifolds, has measure zero in $M$. Smale [Sma65] proved a version of this result in the case both $M$ and $N$ are infinite dimensional Banach manifolds and the map $f$ is Fredholm: in this case the conclusion is that the set of regular values of $f$ is residual in $N$.
When $N$ is infinite-dimensional and $M$ is finite dimensional, $f$ cannot be Fredholm and Smale’s result cannot be applied. In fact there are smooth surjective maps without regular points from any infinite-dimensional Banach space to $\mathbb{R}^{2}$, see [Bat93]. Even under the hypothesis that $f$ is a polynomial map, the set of its critical values can be of positive measure: Kupka [Kup65] constructed an example of a smooth map $f:\ell^{2}\to\mathbb{R}$, whose restriction to each finite dimensional space is a polynomial of degree $3$, and with the property that the set of its critical values is the segment $[0,1]$. On the other hand, using a quantitative version of the classical Sard theorem [Yom83], Yomdin has shown that if a map $f:\ell^{2}\to\mathbb{R}$ can be approximated well-enough with finite dimensional polynomials, still the set of its critical values has measure zero, see [Yom88]. (Kupka’s map does not have this property.) In Yomdin’s proof, the fact that the codomain is one-dimensional is essential and new technical difficulties arise for maps with values in $\mathbb{R}^{m}$, $m>1$.
In this paper we investigate the Sard property for polynomial maps defined on Hilbert spaces and with values in $\mathbb{R}^{m}$, for general $m\geq 1$.
Definition 1.
Given a Hilbert space $H$ and $d,m\in\mathbb{N}$, we define the class of maps
(1) $\mathscr{P}_{d}^{m}(H):=\bigg{\{}f:H\to\mathbb{R}^{m}\,\bigg{|}\,\dim(E)<\infty\implies f|_{E}\textrm{ is a polynomial map of degree $d$}\bigg{\}},$
where by polynomial map we mean that each component of $f|_{E}$ is a polynomial of degree $d$. We require that each element of $\mathscr{P}_{d}^{m}(H)$ is of class $\mathcal{C}^{1}$ (in the Fréchet sense), and that its differential $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous and locally Lipschitz.
The map constructed by Kupka belongs to this family, as well as the Endpoint maps for horizontal path spaces on Carnot groups (see Section 1.3).
We need to introduce a notion of quantitative compactness for subsets of Hilbert spaces. More precisely, given $K\subset H$ and $n\in\mathbb{N}$, we denote by $\Omega_{n}(K,H)$ its Kolmogorov $n$–width:
(2) $\Omega_{n}(K,H):=\inf_{\dim(E)=n}\sup_{u\in K}\inf_{v\in E}\|v-u\|,$
[…]
Our results are stated in terms of entropy dimension of the set of $\nu$-critical values of a map $f:H\to\mathbb{R}^{m}$, namely $f(\operatorname{Crit}_{\nu}(f))$, where $\operatorname{Crit}_{\nu}(f)$ is the set of points where $Df$ has rank at most $\nu\leq m-1$. The classical set of critical values corresponds to setting $\nu=m-1$, and it is denoted by $f(\operatorname{Crit}(f))$, where $\operatorname{Crit}(f)$ is the set of critical points. We refer to Definition 24 of entropy dimension (denoted by $\dim_{e}$, and also known as Minkowski dimension), noting that it is larger than the Hausdorff one so that in the forthcoming estimates one can replace the former with the latter for simplicity.
Our first theorem is a sufficient condition for the validity of the Sard property (see Theorem 31). We denote by $\mu$ the Lebesgue measure on $\mathbb{R}^{m}$.
Theorem A (Sard under $n$–width assumptions).
Let $H$ be a Hilbert space and $K\subset H$ be a compact set such that, for some $q>1$, it holds
(3) $\limsup_{n\to\infty}\Omega_{n}(K,H)^{1/n}\leq q^{-1}.$
Let $d,m\in\mathbb{N}$. There exists $\beta_{0}=\beta_{0}(d,m)>1$ such that for all $f\in\mathscr{P}_{d}^{m}(H)$ and $\nu\leq m-1$ we have
(4) $\dim_{e}\bigg{(}f\big{(}\operatorname{Crit}_{\nu}(f)\cap K\big{)}\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q}.$
In particular, if $q>\beta_{0}$, then the Sard property holds on $K$:
(5) $\mu\bigg{(}f\big{(}\operatorname{Crit}(f)\cap K\big{)}\bigg{)}=0.$
Remark 2.
The constant $\beta_{0}(d,m)$ of Theorem A comes from semialgebraic geometry, and its origin is discussed in Section 1.2. For this reason, we sometimes call it the semialgebraic constant.
By building upon Kupka’s counterexamples [Kup65], we prove that the quantitative assumption $q>\beta_{0}$ in Theorem A cannot be dispensed of. The following result corresponds to Theorem 35. We denote by $B_{X}(r)$ the closed ball of radius $r$ in a Banach space $X$.
Theorem B (Counterexamples to Sard).
Let $d,m\in\mathbb{N}$, with $d\geq 3$, and let $1<q<(d-1)^{1/d}$. There exist a Hilbert space $H$, and $f\in\mathscr{P}_{d}^{m}(H)$ such that $K=\operatorname{Crit}(f)\cap B_{H}(r)$ is compact for all $r>0$, with
(6) $\limsup_{n\to\infty}\Omega_{n}(K,H)^{1/n}\leq q^{-1},$
and $f:H\to\mathbb{R}^{m}$ does not verify the Sard property, namely $\mu(f(\operatorname{Crit}(f)\cap K))>0$. Therefore, the semialgebraic constant $\beta_{0}(d,m)$ of Theorem A satisfies
(7) $\beta_{0}(d,m)\geq(d-1)^{1/d},\qquad\forall\,m\in\mathbb{N},\,d\geq 3.$
Remark 3.
The case $d=2$, which is left out from Theorem B, is special. This is related to the fact that the complexity of semialgebraic sets defined by quadratic equations has a different behaviour compared to the case $d\geq 3$. In this case we currently expect that any $f\in\mathscr{P}^{m}_{2}(H)$ satisfies the Sard property globally on $H$. This will be the subject of a future work.
Replacing the entropy dimension with the Hausdorff dimension (which is smaller than the entropy dimension, but countably stable) we strengthen Theorem A as follows (see Theorem 37).
Theorem C (Sard on linear subspaces).
Let $H$ be a Hilbert space and $K\subset H$ be a compact set such that, for some $q>1$, it holds
(8) $\limsup_{n\to\infty}\Omega_{n}(K,H)^{1/n}\leq q^{-1}.$
Let $d,m\in\mathbb{N}$, let $\beta_{0}=\beta_{0}(d,m)>1$ be the same constant of Theorem A. Consider the (possibly non-closed) linear subspace
(9) $V:=\mathrm{span}(K).$
Then for all $f\in\mathscr{P}_{d}^{m}(H)$ and $\nu\leq m-1$ the restriction $f|_{V}:V\to\mathbb{R}^{m}$ satisfies
(10) $\dim_{\mathcal{H}}\bigg{(}f\big{(}\mathrm{Crit}_{\nu}(f|_{V})\big{)}\bigg{)}\leq\nu+\frac{\log\beta_{0}}{\log q}.$
In particular, if $q>\beta_{0}$, then the restriction $f|_{V}:V\to\mathbb{R}^{m}$ satisfies the Sard property
(11) $\mu\bigg{(}f\big{(}\mathrm{Crit}(f|_{V})\big{)}\bigg{)}=0.$
In all previous results, $f\big{(}\mathrm{Crit}(f|_{V})\big{)}$ can be replaced with the smaller set $f(\mathrm{Crit}(f)\cap V)$.
Remark 4.
The restriction $f|_{V}:V\to\mathbb{R}^{m}$ is a $\mathcal{C}^{1}$ map in the Fréchet sense from the normed vector space $(V,\|\cdot\|)$ to $\mathbb{R}^{m}$, with $D_{u}(f|_{V})=(D_{u}f)|_{V}$ for all $u\in V$.
Example 5.
Let $q\in(1,\infty)$ with $q>\beta_{0}(d,m)$, where $\beta_{0}(d,m)>1$ is the number in Theorems A and C, for $d,m\in\mathbb{N}$. Let $H=\ell^{2}$ with the usual norm. Consider the subset $K\subset B_{H}(1)$ given by
(12) $K=\left\{u\in\ell^{2}\,\middle|\,\sum_{j=1}^{\infty}|u_{j}|^{2}q^{2j}\leq 1\right\}.$
One can easily verify that $K$ is compact. Furthermore, the linear space $V=\mathrm{span}(K)$ is dense in $H$. We can estimate the $n$–width of $K$ as follows. For $n\in\mathbb{N}$ consider the $n$-dimensional subspaces $E_{n}=\{x\in\ell^{2}\mid x_{j}=0,\,\forall\,j\geq n+1\}\subset H$. From (2), we obtain
(13) $\Omega_{n}(K,H)^{2}$ $\leq\sup_{u\in K}\inf_{v\in E_{n}}\|u-v\|^{2}$
(14) $=\sup_{u\in K}\sum_{j=n+1}^{\infty}|u_{j}|^{2}$
(15) $\leq q^{-2(n+1)}\sup_{u\in K}\sum_{j=1}^{\infty}|u_{j}|^{2}q^{2j}\leq q^{-2(n+1)}.$
Hence $\limsup_{n}\Omega_{n}(K,H)^{1/n}\leq q^{-1}$, with $q>\beta_{0}(d,m)$. Thus, by Theorem C it holds (in particular)
(16) $\mu\bigg{(}f\big{(}\mathrm{Crit}(f|_{V})\big{)}\bigg{)}=\mu\bigg{(}f\big{(}\mathrm{Crit}(f)\cap V\big{)}\bigg{)}=0,\qquad\forall\,f\in\mathscr{P}_{d}^{m}(H).$
Note that even if $V\subset H$ is dense, the unrestricted map $f:H\to\mathbb{R}^{m}$ may not have the Sard property (see e.g. the Kupka counterexamples in Section 4.3).
Theorems A, F and C are deduced from a more general result for maps that are “well-approximated” by polynomials. This can be regarded as our main result concerning the Sard property, and does not make use of the concept of $n$–width. We report here the statement (see Theorem 25).
Theorem D (Sard criterion for well-approximated maps).
Let $d,m\in\mathbb{N}$. There exists a constant $\beta_{0}=\beta_{0}(d,m)>1$ such that the following holds. Let $H$ be a Hilbert space, and let $f:H\to\mathbb{R}^{m}$ be a $\mathcal{C}^{1}$ map such that its differential $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous. Let $K\subset H$ be a bounded set with this approximation property: there exist a sequence $E_{n}\subset H$ of linear subspaces, $\dim(E_{n})=n$, and polynomial maps $f_{n}:E_{n}\to\mathbb{R}^{m}$ with uniformly bounded degree:
(17) $\sup_{n\in\mathbb{N}}\deg f_{n}\leq d<\infty,$
such that for some $q>1$, $c\geq 0$, and all large enough $n$ it holds
(18) $\sup_{x\in K}\bigg{(}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|+\|\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}\bigg{)}\leq cq^{-n}.$
Then for all $\nu\leq m-1$ it holds
(19) $\dim_{e}\bigg{(}f\big{(}\operatorname{Crit}_{\nu}(f)\cap K\big{)}\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q}.$
In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property on $K$:
(20) $\mu\bigg{(}f(\operatorname{Crit}(f)\cap K)\bigg{)}=0.$
By means of the general Theorem D we single out a class of maps satisfying the Sard property on the whole domain of definition. This is the content of Theorem 41, that we report here. The case $m=1$ corresponds to [YC04, Thm. 10.12].
Theorem E (Global Sard for special maps).
Let $H$ be a separable Hilbert space. For all $k\in\mathbb{N}$, let $p_{k}:E_{k}\to\mathbb{R}^{m}$ be polynomial maps with $\sup_{k\in\mathbb{N}}\deg p_{k}\leq d$ for some $d\in\mathbb{N}$, and such that
(21) $\sup_{x\in B_{E_{k}}(1)}\|p_{k}(x)\|\leq q^{-k},\qquad\forall\,k\in\mathbb{N},$
for some $q>1$. Then the map $f:H\to\mathbb{R}^{m}$ defined by
(22) $f(x):=\sum_{k=1}^{\infty}p_{k}(x_{1},\dots,x_{k}),\qquad\forall\,x\in H,$
is well–defined, $f\in\mathscr{P}_{d}^{m}(H)$ (see Definition 1), and for all $\nu\leq m-1$ and $r>0$ it holds
(23) $\dim_{e}\bigg{(}f\left(\operatorname{Crit}_{\nu}(f)\cap B_{H}(r)\right)\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q},$
where $\beta_{0}=\beta_{0}(d,m)>1$ is the same constant given by Theorem D. In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property globally on $H$:
(24) $\mu\bigg{(}f\big{(}\operatorname{Crit}(f)\big{)}\bigg{)}=0.$
As a consequence of Theorems A and B, we see that there is a threshold on the $n$–width for the validity of the Sard property. Motivated by this fact, for a compact set $K\subset H$ we define the quantity:
(25) $\omega(K,H):=\limsup_{n\to\infty}\Omega_{n}(K,H)^{1/n}\in[0,1].$
Roughly speaking, smaller $\omega(K,H)$ means that $K$ is “more compact” in $H$, and better approximated by finite-dimensional subspaces. We obtain the following statement (see Theorem 42).
Theorem F (Sard threshold on compacts).
For all $d,m\in\mathbb{N}$, there exists $\omega_{0}(d,m)\in(0,1]$ such that
  1. (i)
    for every $f\in\mathscr{P}_{d}^{m}(H)$ and for every compact set $K\subset H$ with $\omega(K,H)<\omega_{0}(d,m)$,
    (26) $\mu\bigg{(}f\left(\mathrm{crit}(f)\cap K\right)\bigg{)}=0;$
  2. (ii)
    for every $\omega>\omega_{0}(d,m)$, with $\omega\in(0,1]$, there exist $f\in\mathscr{P}_{d}^{m}(H)$ and a compact set $K\subset H$ with $\omega(K,H)=\omega$ and such that
    (27) $\mu\bigg{(}f\left(\mathrm{crit}(f)\cap K\right)\bigg{)}>0.$
Remark 6.
By construction, and Theorem A, $\omega_{0}\geq\beta_{0}^{-1}$. Furthermore, if $d\geq 3$, Theorem B yields $\omega_{0}<1$, so that Item ii is non-vacuous in these cases.
1.2. Quantitative semialgebraic geometry
The proof of Theorem D needs some fine properties of semialgebraic sets. The strategy of the proof is conceptually similar to Yomdin’s [Yom88], and uses the theory of variations, introduced by Vitushkin [Vit55, Vit61] and developed in [YC04].
Let us denote with $V_{i}(S)$ the $i$-th variation of a semialgebraic set, which is a sort of $i$-dimensional volume, (see Definition 14). Furthermore, recall that for a $C^{1}$ map $F:\mathbb{R}^{n}\to\mathbb{R}^{m}$, and $\Lambda=(\Lambda_{1},\ldots,\Lambda_{m})\in\mathbb{R}^{m}_{+}$ the set of almost-critical values of $F$ is defined by
(28) $C^{\Lambda}(F):=\bigg{\{}x\in\mathbb{R}^{n}\,\bigg{|}\,\sigma_{i}(D_{x}F)\leq\Lambda_{i},\quad\forall\,i=1,\dots,m\bigg{\}},$
where $\sigma_{1}(D_{x}F)\geq\cdots\geq\sigma_{m}(D_{x}F)$ are the singular values of $D_{x}F$, see Definition 16 (here we assume $n\geq m$). In [YC04, Cor. 7.4] a quantitative estimate on the variations of the almost-critical values of polynomial maps has been obtained. In that estimate the dimensional parameter $n$ does not appear explicitly. We obtain Theorem G below, which makes this dependence explicit (see Theorem 23).
Theorem G (Quantitative variations estimates).
Let $n\geq m$, and $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a polynomial map with components of degree at most $d$. For $i=0,\dots,m$, $\Lambda=(\Lambda_{1},\dots,\Lambda_{m})\in\mathbb{R}^{m}_{+}$ and $r>0$, we have
(29) $V_{i}(p(C^{\Lambda}(p)\cap B_{\mathbb{R}^{n}}(r)))\leq\mathrm{cst}(m,r)n^{m}\beta_{0}^{n}\Lambda_{0}\cdots\Lambda_{i},$
where $\beta_{0}=\beta_{0}(d,m)>1$ depends only on $d$ and $m$, $\mathrm{cst}(m,r)$ depends only on $m,r$, and we set $\Lambda_{0}=1$.
A key technical step in the proof of Theorem G is a quantitative substitute for the definable choice theorem in semialgebraic geometry, which we prove in [LRT24] (see Theorem 13).
1.3. Applications to Endpoint maps of Carnot groups
We discuss now the implications of Theorem F and Theorem C in the context of sub-Riemannian geometry.
Recall that an $m$-dimensional Carnot group of step $s\in\mathbb{N}$ is a connected and simply connected Lie group $(\mathbb{G},\cdot)$ of dimension $m$, whose Lie algebra $\mathfrak{g}$ admits a stratification of step $s$, that is
(30) $\mathfrak{g}=\mathfrak{g}_{1}\oplus\dots\oplus\mathfrak{g}_{s},$
where $\mathfrak{g}_{i}\neq\{0\}$, $\mathfrak{g}_{i+1}=[\mathfrak{g}_{1},\mathfrak{g}_{i}]$ for all $i=1,\dots,s-1$ and $[\mathfrak{g}_{1},\mathfrak{g}_{s}]=\{0\}$. The group exponential map yields an identification $\mathbb{G}\simeq\mathbb{R}^{m}$, and the first stratum $\mathfrak{g}_{1}$ of the Lie algebra defines a smooth, totally non-holonomic distribution $\Delta\subseteq T\mathbb{R}^{m}$ of rank $k:=\dim\mathfrak{g}_{1}$.
Fix a global trivialization of $\Delta$:
(31) $\Delta=\mathrm{span}\{X_{1},\ldots,X_{k}\},$
where each $X_{i}$ is a left-invariant vector field. Let $I:=[0,1]$ be the unit interval and $H:=L^{2}(I,\mathbb{R}^{k})$. We say that a curve $\gamma:I\to\mathbb{G}$ is horizontal if it is absolutely continuous and there exist $u\in H$, called control such that for a.e. $t\in I$ it holds
(32) $\dot{\gamma}(t)=\sum_{i=1}^{k}u_{i}(t)X_{i}(\gamma(t)),$
Furthermore, for any given $u\in H$ there exists a unique $\gamma_{u}:I\to\mathbb{G}$ satisfying (32) and such that $\gamma_{u}(0)=e$ (the identity $e\in\mathbb{G}$, identified with $0\in\mathbb{R}^{m}$). Note that the class of horizontal curves does not depend on the choice of the global trivialization of $\Delta$.
The Endpoint map is the map that sends a control $u$ to the the corresponding final point $\gamma_{u}(1)$:
(33) $\mathcal{E}:L^{2}(I,\mathbb{R}^{k})\to\mathbb{G}.$
The Sard conjecture is the conjecture that Endpoint map has the Sard property:
Sard conjecture: the set $\mathcal{E}(\mathrm{Crit}(\mathcal{E}))$ has zero measure.
The conjecture was introduced by Zhitomirskii and Montgomery, see [Mon02, Sec. 10.2] for general sub-Riemannian structures, where the Endpoint map can be defined in a similar way, on a suitable domain of horizontal paths, which has the structure of a Hilbert manifold. For the specific case of Carnot groups, it is known to be true for step $\leq 2$, see [AGL15, LDMO+16]. Furthermore, the conjecture has been verified in [BNV22] for filiform Carnot groups, in [BV20] for Carnot groups of rank 2 and step 4, and rank 3 and step 3, and for a handful of other specific examples described in [LDMO+16].
Assume that $\Delta$ is also endowed with a left-invariant norm. This choice induces a length structure on the space of horizontal curves, and correspondingly, a distance, which gives $\mathbb{G}$ the structure of sub-Finsler Carnot group (or sub-Riemannian Carnot group if the norm is induced by a scalar product). In this case, within the set critical points of $\mathcal{E}$ (also called singular horizontal paths), the energy-minimizing ones play a significant role. At the corresponding critical values, the distance loses regularity: it is never smooth and it can even lose local semiconcavity [BR19, Sec. 4.2]. For this reason, the understanding of the “abundance” of such critical values is crucial. Denoting with $H_{\min}\subseteq L^{2}(I,\mathbb{R}^{k})$ the set of minimizing horizontal paths starting from the identity, the minimizing Sard conjecture can be formulated as follows:
Minimizing Sard conjecture: the set $\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap H_{\min})$ has zero measure.
The most general result to date is that the above set is a closed nowhere dense set [Agr09, RT05], which of course does not imply the minimizing Sard conjecture. The above conjecture is one of the main open problems in sub-Riemannian geometry [Agr14, Prob. 3], [RT05, Conj. 1]. In general, the problem is settled in the following cases:
    Thanks to the work of Agrachev-Sarychev [AS96] and Agrachev-Lee [AL09], in absence of so-called Goh singular minimizing paths, all sub-Riemannian minimizing paths (resp. sub-Finsler, under suitable smoothness assumptions on the norm) are solution of a finite-dimensional Hamiltonian flow so that the corresponding set of critical values has zero measure by the finite-dimensional Sard theorem. Recently, Rifford proved that, more in general, the minimizing Sard conjecture holds for structures where all non-trivial Goh paths have Goh-rank $\leq 1$ almost everywhere, see [Rif23]. For Carnot groups, these results can be applied for example when the distribution is pre-medium fat, see [Rif23, (1.2)]. In passing, we remark that for generic (so, typically not Carnot) sub-Riemannian structures with distribution of rank $\geq 3$, there are no non-trivial singular Goh minimizing paths, see [CJT06, Cor. 2.5] and [AG01, Thm. 8]. As a consequence the minimizing Sard conjecture holds true. Thanks to [Rif23, Cor. 1.4], the latter result extends the case of rank $\geq 2$.
    The minimizing Sard conjecture is true also for Carnot groups of step $\leq 3$, see [LDMO+16]. In this case, the problem is reduced (in a non-trivial way) to a finite-dimensional one. First, by noting that singular minimizing curves are non-singular in some proper subgroup, and then exploiting the fact that Carnot subgroups are parametrized by a finite-dimension manifold.
Finally, we refer to [ZZ95, BdSR18, BdSFPR22, BdSPR23, BdSPR22, OV19, RT05] for further details and discussions of various forms of the Sard conjecture on general sub-Riemannian structures.
The following result (Theorem 46) connects this framework to the previous sections.
Theorem H (Polynomial properties of the Endpoint map).
Let $\mathbb{G}$ be a Carnot group of topological dimension $m$, step $s$, and rank $k$. Then the Endpoint map $\mathcal{E}\in\mathscr{P}_{s}^{m}(H)$, for $H=L^{2}(I,\mathbb{R}^{k})$
As a consequence of Theorem H, we can apply Theorem C to the study of Endpoint maps of Carnot groups. We record here a first immediate corollary.
Theorem I (Sard criterion for Endpoint maps).
Let $\mathbb{G}$ be a Carnot group of topological dimension $m$, step $s$, and rank $k$. Let $K\subset L^{2}(I,\mathbb{R}^{k})$ be a compact subset and let $q>1$ be such that
(34) $\limsup_{n\to\infty}\Omega_{n}(K,L^{2}(I,\mathbb{R}^{k}))^{1/n}\leq q^{-1}.$
Then, letting $V:=\mathrm{span}(K)$, the restriction $\mathcal{E}|_{V}:V\to\mathbb{G}$ satisfies
(35) $\dim_{\mathcal{H}}\bigg{(}\mathcal{E}\big{(}\mathrm{Crit}(\mathcal{E}|_{V})\big{)}\bigg{)}\leq m-1+\frac{\log\beta_{0}}{\log q},$
where $\beta_{0}=\beta_{0}(s,m)>1$ is the same constant of Theorem A. In particular, it holds:
  1. (i)
    if $q>\beta_{0}$, the restriction $\mathcal{E}|_{V}:V\to\mathbb{G}$ has the Sard property, namely
    (36) $\mu\bigg{(}\mathcal{E}\big{(}\mathrm{Crit}(\mathcal{E}|_{V})\big{)}\bigg{)}=0.$
  2. (ii)
    if $q=\infty$, the restriction $\mathcal{E}|_{V}:V\to\mathbb{G}$ has the so-called strong Sard property, namely
    (37) $\dim_{\mathcal{H}}\bigg{(}\mathcal{E}\big{(}\mathrm{Crit}(\mathcal{E}|_{V})\big{)}\bigg{)}\leq m-1.$
In all previous results, $\mathcal{E}\big{(}\mathrm{Crit}(\mathcal{E}|_{V})\big{)}$ can be replaced with the smaller set $\mathcal{E}\big{(}\mathrm{Crit}(\mathcal{E})\cap V\big{)}$.
We can apply Theorem I to find large sets on which the Sard property holds for the Endpoint map of Carnot groups. This is a more functional-analytic approach to the Sard conjecture that, unlike previous ones, does not resort to reduction to finite-dimensional cases.
An interesting case in which the assumptions of Theorem I can be effectively checked is given by sets of controls that are “sufficiently regular”. We introduce the notation to state our result, see Section 5.2 for additional details. Given $r>0$, and a closed interval $I\subset\mathbb{R}$, consider the set:
(38) $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r):=\bigg{\{}u:I\to\mathbb{R}^{k}\,\bigg{|}\,\textrm{$u$ is real-analytic with radius of convergence}>r\bigg{\}},$
endowed with the sup norm over the $r$-neighbourhood of $I$ in $\mathbb{C}$. More in general, let $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ be the set of piecewise analytic controls:
(39) $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)=\bigg{\{}u:I\to\mathbb{R}^{k}\,\bigg{|}\,u|_{I_{j}}\in\mathcal{C}^{\omega}(I_{j},\mathbb{R}^{k};r),\,\text{ for all $j=1,\dots,\ell$}\bigg{\}},$
where $I_{j}=\inf I+\left[\frac{(j-1)|I|}{\ell},\frac{j|I|}{\ell}\right]$, for all $j=1,\dots,\ell$, which we equip with a suitable norm (see Section 5.2). In Theorem 47 we prove that the unit ball $K$ of $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$, with $I=[0,1]$, is compact in $H$ and its $n$–width satisfies:
(40) $\Omega_{n}(K,H)\leq\frac{(k\ell)^{1/2}}{\ln r}\left(\frac{1}{r}\right)^{\left\lfloor\frac{n}{k\ell}\right\rfloor},\qquad\forall\,n\in\mathbb{N},\quad r>1.$
By applying Theorem I we can therefore deduce the following result (see Theorem 48).
Theorem J (Sard property on piecewise real–analytic controls).
Let $\mathbb{G}$ be a Carnot group of topological dimension $m$, step $s$, and rank $k$. Given $\ell\in\mathbb{N}$, there exists $r=r(m,s,k,\ell)>0$ such that, letting $V=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$, with $I=[0,1]$, it holds
(41) $\mu\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap V)\big{)}=\mu\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{V})\big{)}=0.$
Namely, the Sard property holds on the space of piecewise real–analytic controls with radius of convergence $>r$ and with $\ell$ pieces. Furthermore, letting $W=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};\infty,\ell)$, then it holds
(42) $\dim_{\mathcal{H}}\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap W)\big{)}\leq\dim_{\mathcal{H}}\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{W})\big{)}\leq m-1.$
Namely, the strong Sard property holds on the space of piecewise entire controls with $\ell$ pieces.
Remark 7 (Regularity and Sard).
Sussmann established that for real–analytic sub-Riemannian structures (such as Carnot groups), minimizing horizontal paths are real–analytic on an open and dense set of their interval of definition [Sus14]. In light of that result, Theorem J hints at a unexpected link between the regularity problem of sub-Riemannian geodesics and the minimizing Sard conjecture.
Remark 8 (Surjectivity of the Endpoint map).
Of course a subset of controls $V\subset L^{2}(I,\mathbb{R}^{k})$ must satisfy $\mathcal{E}(V)=\mathbb{G}$ to be relevant for the theory. By the Rashevskii-Chow theorem, $\mathcal{E}:L^{2}(I,\mathbb{R}^{k})\to\mathbb{R}^{m}$ is surjective. We observe that $\mathcal{E}$ is surjective also on $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ if $\ell\geq 4m$, since from the proof of the Rashevskii-Chow theorem any pair of points can be joined by a concatenation of $4m$ horizontal paths with constant controls, see [ABB20, Sec. 3.2]. Actually, the Endpoint map is surjective when restricted on any set that is dense in $L^{2}(I,\mathbb{R}^{k})$, see e.g. [Bel96, Sec. 2.5]. It follows that $\mathcal{E}$ is surjective on the sets of Theorem J since in any case they contain all (restrictions on $I$ of) polynomial controls.
Finally, we prove in Theorem 52 that the Endpoint map on Carnot groups is surjective already when restricted to polynomial controls of large enough degree (depending on $\mathbb{G}$). This is a particular case of the following more general result (see Theorem 53).
Theorem K (Quantitative surjectivity).
Let $\mathbb{G}$ be a Carnot group with topological dimension $m$ and rank $k$. Let $S\subset L^{2}(I,\mathbb{R}^{k})$ be a dense set. Then there exist $u_{0},u_{1},\dots,u_{m}\in S$ such that
(43) $\mathcal{E}(\mathrm{span}\{u_{0},u_{1},\dots,u_{m}\})=\mathbb{G}.$
In other words, one can find an $(m+1)$-dimensional vector subspace $E\subset L^{2}(I,\mathbb{R}^{k})$ of controls (depending on $\mathbb{G}$) such that the restriction of the Endpoint map to $E$ is surjective. Furthermore, $E$ can be assumed to have generators in any prescribed dense set $S\subset L^{2}(I,\mathbb{R}^{k})$. Theorem K can be seen as a more refined version (for Carnot groups) of the well–known fact that the Endpoint map is surjective when restricted on any set of controls that is dense in $L^{2}(I,\mathbb{R}^{k})$, see e.g. [Bel96, Sec. 2.5].
1.4. Closing thoughts on the Sard conjecture
We do not venture in a guess in favor of Sard conjecture on the whole $H=L^{2}(I,\mathbb{R}^{k})$, instead we propose a direction of future investigation that can be approached with our techniques. Recall that $H_{\mathrm{min}}\subset H$ is the set of energy-minimizing controls. It is known that $H_{\min}$ is boundedly compact in $H$ i.e., the intersection $H_{\min}\cap B_{H}(r)$ is compact for all $r>0$, see [ABB20, Thm. 8.66], [Agr98]. In fact, from the very recent result by Lokutsievskiy and Zelikin [LZ23, Cor. 1 $(\mathfrak{G})$], one can deduce the following estimate for the $n$–width on Carnot groups
(44) $\Omega_{2n+1}(H_{\min}\cap B_{H}(1),H)\leq Cn^{-\frac{1}{2s}},\qquad\forall\,n\in\mathbb{N},$
for some $C>0$ depending on $\mathbb{G}$. Estimate (44) can be understood as a quantitative compactness property for $H_{\min}$. Unfortunately, the polynomial decay (w.r.t. $n$) of (44) is too weak to apply Theorem A, since the latter requires an exponential one.
If one could prove that, for some $q^{-1}<\omega_{0}(s,m)$, it holds
(45) $\Omega_{n}(H_{\min}\cap B_{H}(1),H)\leq q^{-n},\qquad\text{as }n\to\infty,$
then Theorem F would settle the minimizing Sard conjecture on Carnot groups.
We also note that it would be sufficient to prove (45) for the set of the so-called strictly abnormal energy-minimizing controls $H_{\min}^{\mathrm{str.abn}}\subsetneq H_{\min}$. Unfortunately, we were not able to produce a direct estimate of the $n$–width of (bounded subsets of) $H_{\min}$ or $H_{\min}^{\mathrm{str.abn}}$, so that we have no further evidence to support this idea towards a proof of the minimizing Sard conjecture.
In the opposite direction, and to conclude, our results can be used to identify spaces where the (general) Sard conjecture can be violated. If, for a given Carnot group $\mathbb{G}$, there is a linear subset of controls $V\subset H$ such that $\mu(\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap V))>0$, then the following necessary condition must hold:
(46) $\limsup_{n\to\infty}\Omega_{n}(V\cap B_{H}(r),H)^{1/n}\geq\omega_{0}(s,m).$
1.5. Acknowledgements
This project has received funding from (i) the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement GEOSUB, No. 945655); (ii) the PRIN project “Optimal transport: new challenges across analysis and geometry” funded by the Italian Ministry of University and Research; (iii) the Knut and Alice Wallenberg Foundation. The authors also acknowledge the INdAM support. The authors also wish to thank Yosef Yomdin for helpful discussions that improved the presentation of the paper.
2. Table of notations
{xltabular}
p0.19 X +$\mathbb{R}^{m}_{+}$ $m$-tuples of positive real numbers
$B_{X}(x,r)$ closed ball of radius $r>0$ centered at $x$ of a Banach space $X$
$B_{X}(r)$ closed ball of radius $r>0$ centered at 00 of a Banach space $X$
$B_{X}$ closed unit ball centered at 00 of a Banach space $X$
$\mathcal{L}(X,Y)$ linear continuous maps between banach spaces $X$, $Y$
$\|\cdot\|_{\operatorname{op}}$ operator norm on $\mathcal{L}(X,Y)$, see (119)
$\|\cdot\|_{X}$ norm for a Banach space $X$ (the $X$ is omitted when there is no confusion)
$\mathcal{B}(X,Y)$ continuous bounded maps between Banach spaces $X$, $Y$
$\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ piece-wise real–analytic functions, Section 5.2
$\mu$ Lebesgue measure
$\Omega_{n}(K,H)$ Kolmogorov $n$–width of $K\subset H$, Definition 26
$\omega(K,H)$ asymptotic Kolmogorov width of $K\subset H$, Eq. 25
$\mathcal{P}_{d}^{m}(H)$ set of well-approximable maps, Definition 1
$\beta_{0}(d,m)$ semialgebraic constant, Remark 2
$\omega(d,m)$ threshold constant, Theorem F
$\mathcal{H}^{i}(A)$ $i$-dimensional Hausdorff measure of a rectifiable Borel set $A$
$V_{i}(A)$ Vitushkin variation of a bounded semialgebraic set $A$, Definition 14
$b_{0}(S)$ number of connected components of a set $S$
$\sigma_{k}(L)$ $k$-th singular value of a linear map $L:H\to\mathbb{R}^{m}$ on a Hilbert space, ordered in decreasing order $\sigma_{1}(M)\geq\dots\geq\sigma_{m}$, Definition 16
$\operatorname{Crit}(f)$ critical points of a map, Eq. 115
$\operatorname{Crit}_{\nu}(f)$ $\nu$-critical points of a map, Eq. 115
$C^{\Lambda}(f)$ $\Lambda$-critical (or almost-critical) points of a map, Definition 19
$M(\varepsilon,S)$ $\varepsilon$-entropy of a set $S\subset\mathbb{R}^{m}$, Definition 24
$\dim_{e}(S)$ entropy dimension of a set $S\subset\mathbb{R}^{m}$, Definition 24
$\mathcal{U}_{\varepsilon}(S)$ $\varepsilon$-neighbourhood of a subset $S\subset\mathbb{R}^{n}$
$\mathrm{dist}_{H}$ Hausdorff distance
$\beta_{\mathrm{tm}}$ Thom-Milnor bound constant, Theorem 12
$\beta_{\mathrm{dc}}$ definable choice constant, Theorem 13
$\beta_{\mathrm{yc}}$ Yomdin-Comte constant, Theorem 18
$D(S)$ diagram of a semialgebraic set $S$, Definition 10
$\mathbb{G}$ Carnot group (of step $s$, rank $k$, dimension $m$), Definition 43
$\mathcal{E}$ Endpoint map, Definition 45
3. Quantitative variations estimates
3.1. Semialgebraic sets and maps
The goal of this section is to prove Theorem G. We begin by recalling some basic notions from semialgebraic geometry.
Definition 9 (Semialgebraic sets and maps).
We say that a set $S\subset\mathbb{R}^{n}$ is semialgebraic if
(47) $S=\bigcup_{i=1}^{a}\bigcap_{j=1}^{b_{i}}\left\{x\in\mathbb{R}^{n}\mid\mathrm{sign}(p_{ij}(x))=\sigma_{ij}\right\},$
for some finite set of polynomials $p_{ij}\in\mathbb{R}[x_{1},\ldots,x_{n}]$ and $\sigma_{ij}\in\{0,+1,-1\}$, where
(48) $\mathrm{sign}(r):=\begin{cases}+1&r>0,\\ -1&r<0,\\ 0&r=0.\end{cases}$
We call the description (47) a representation of $S$.
A map $f:A\to B$ between semialgebraic sets $A\subset\mathbb{R}^{n},B\subset\mathbb{R}^{\ell}$ is said to be semialgebraic if its graph is a semialgebraic set in $\mathbb{R}^{n}\times\mathbb{R}^{\ell}$.
The representation of a semialgebraic set $S$ as in (47) is not unique. However, a representation is useful to quantify the complexity of $S$, using the following notion.
Definition 10 (Diagram of a semialgebraic set).
Let $S\subset\mathbb{R}^{n}$ be a semialgebraic set represented as in (47). We will say that the triple
(49) $\left(n,a\cdot\max_{i}\{b_{i}\},\max_{i,j}\{\deg(p_{ij})\}\right)\in\mathbb{N}^{3}$
is a diagram for $S$. Below, the equation “$D(S)=(m,c,d)$” will mean that there exists a representation of $S$ as in (47) with $n\leq m$, $a\cdot\max_{i}\{{b_{i}}\}\leq c$ and $\max_{i,j}\{\deg(p_{ij})\}\leq d$.
Remark 11.
There are several ways of quantifying the complexity of a semialgebraic set, depending on the way it is presented. Essentially, these notions should contain three pieces of information: the number of variables, the “combinatorics” of the presentation, and the degrees of the defining polynomials. For instance, every semialgebraic set $S\subset\mathbb{R}^{n}$ can be presented as
(50) $S=\bigcup_{i=1}^{a}\bigcap_{j=1}^{b_{i}}E_{ij},$
where each $E_{ij}$ has the form $\{x\in\mathbb{R}^{n}\mid p_{ij}(x)\leq 0\}$ or $\{x\in\mathbb{R}^{n}\mid p_{ij}(x)<0\}$, where $p_{ij}$ is a polynomial with $\deg(p_{ij})\leq d$ for all $i,j$. Given a presentation of this type, it is immediate to see that $S$ also admits a diagram $D(S)=(n,c,d)$ (i.e. it can be presented as in (47) with $c=c(a,b_{1},\dots,b_{a})$.
3.1.1. Dimension and stratifications
Every semialgebraic subset of $\mathbb{R}^{n}$ can be written as a finite union of semialgebraic sets, each of them semialgebraically homeomorphic to an open cube $(0,1)^{m}\subset\mathbb{R}^{m}$, for some $m\leq n$ ([BCR98, Thm. 2.3.6]). This allows to define the dimension of a semialgebraic set as the maximum of the dimensions $m$ of these cubes. Every semialgebraic set can be written as a finite union of smooth, semialgebraic disjoint manifolds called strata ([BCR98, Prop. 9.1.8]).
Notice that the dimension of a semialgebraic set is preserved by semialgebraic homeomorphisms and behaves naturally under product structures:
(51) $\dim(A\times B)=\dim(A)+\dim(B).$
Moreover, if $f:A\to B$ is a continuous semialgebraic map, then
(52) $\dim(f(A))\leq\dim(A).$
3.1.2. Semialgebraic triviality
Continuous semialgebraic maps $f:A\to B$ are “piecewise” trivial fibrations: there exists a partition of $B$ into finitely many semialgebraic sets
(53) $B=\bigsqcup_{j=1}^{b}B_{j}$
and, for every $j=1,\ldots,b$ there exist fibers $F_{j}:=f^{-1}(b_{j})$, with $b_{j}\in B_{j}$ and a semialgebraic homeomorphism $\psi_{j}:f^{-1}(B_{j})\to B_{j}\times F_{j}$ that make the following diagram commutative:
(54) ${B_{j}\times F_{j}}$${f^{-1}(B_{j})}$${B_{j}}$$\scriptstyle{\psi_{j}}$$\scriptstyle{p_{1}}$$\scriptstyle{f}$
This result is called semialgebraic triviality, see [BPR06, Thm. 5.46].
3.1.3. Projections
The image of a semialgebraic set $S\subset\mathbb{R}^{n}\times\mathbb{R}^{\ell}$ under the projection map $\pi:\mathbb{R}^{n}\times\mathbb{R}^{\ell}\to\mathbb{R}^{n}$ is semialgebraic ([BCR98, Thm. 2.2.1]). Moreover, it follows from [BPR06, Thm. 14.16, Notation 8.2] [^1] Note that the role of $k$ and $\ell$ for us is swapped with respect to [BPR06, Thm. 14.16]: we eliminate $\ell$ variables from a list of $n+\ell$, they eliminate $k$ variables from a list of $k+\ell$. that there exists $C_{1}>0$ such that if $(n+\ell,c,d)$ is a diagram for $S$, we can write
(55) $\pi(S)=\bigcup_{i\in I}\bigcap_{j\in J_{i}}\bigcup_{k\in N_{ij}}\left\{x\in\mathbb{R}^{n}\mid\mathrm{sign}(p_{ijk}(x))=\sigma_{ijk}\right\},$
with
(56) #$\#I\leq c^{(n+1)(\ell+1)}d^{C_{1}\ell n},\quad\#J_{i}\leq c^{\ell+1}d^{C_{1}\ell},\quad\#N_{ij}\leq d^{C_{1}\ell},$
and with
(57) $\deg(p_{ijk})\leq d^{C_{1}\ell}.$
In particular, there exists $C_{2}>0$ such $\pi(S)$ admits a diagram with:
(58) $D(\pi(S))=\left(n,(cd)^{C_{2}\ell n}(d^{C_{2}\ell})^{(cd)^{C_{2}\ell}},d^{C_{2}\ell}\right).$
In particular, the diagram of the projection of a semialgebraic set is controlled explicitly by the diagram of the original set.
3.1.4. Thom-Milnor bound
We will need the following qualitative bound on the number of connected components of a semialgebraic set, that follows from [BPR06, Thm. 7.50].
Theorem 12 (Thom-Milnor).
Let $c,d\in\mathbb{N}$. There exists a constant $\beta_{\mathrm{tm}}=\beta_{\mathrm{tm}}(c,d)>1$, such that for any semialgebraic set $S\subset\mathbb{R}^{n}$ with diagram $D(S)=(n,c,d)$ it holds
(59) $b_{0}(S)\leq\beta_{\mathrm{tm}}^{n},$
where $b_{0}(S)$ denotes the number of connected components of $S$.
3.2. Definable choice
We recall from [LRT24] the following quantitative approximate definable choice result, which will be a key tool in the proof of Theorem G.
Let us explain the name and the context of this result. Given a polynomial map $F:\mathbb{R}^{n}\to\mathbb{R}^{\ell}$ and a compact semialgebraic set $A\subset\mathbb{R}^{n}$, for every $y\in F(A)$ one would like to choose a point $x(y)\in F^{-1}(y)\cap A$ so that the resulting set $C:=\{x(y)\}_{y\in F(A)}$ is semialgebraic and with controlled dimension. In particular, the “choice” of such $x(y)$ has to be “definable”. In this way $C$ would be of dimension $\ell$. The fact that this can be done follows from semialgebraic triviality (see Section 3.1.2). However, in this way there is no control on the geometry of $C$ (in particular, the number of its connected components) in terms of the data defining $A$ (namely, its diagram, see Definition 10).
For many geometric applications (as in this paper) one does not really need that $C$ is a choice over $A$, but it is enough that it is close to it. More precisely, given $\varepsilon>0$, denote by $\mathcal{U}_{\varepsilon}(A)$ the Euclidean $\varepsilon$–neighbourhood of $A$. Then one can relax the requirements for the definable choice and ask, given $\varepsilon>0$, for a set $C_{\varepsilon}\subset\mathcal{U}_{\varepsilon}(A)$, of dimension $\leq\ell$, whose image under $F$ is $\varepsilon$-close to $F(A)$ (in the Hausdorff metric, denoted by $\mathrm{dist}_{H}$), and with a control the geometry of $C_{\varepsilon}$.
Theorem 13 (Quantitative approximate definable choice [LRT24]).
For every $c,d,\ell\in\mathbb{N}$ there exists $\beta_{\mathrm{dc}}>1$ satisfying the following statement. Let $n\in\mathbb{N}$ and let $A\subset\mathbb{R}^{n}$ be a closed semialgebraic set contained in the ball $B_{\mathbb{R}^{n}}(r)$ with diagram
(60) $D(A)=(n,c,d).$
Let also $F=(F_{1},\ldots,F_{\ell}):\mathbb{R}^{n}\to\mathbb{R}^{\ell}$ be a polynomial map with components of degree bounded by $d$. Then for every $\varepsilon\in(0,r)$ there exists a closed semialgebraic set $C_{\varepsilon}\subset\mathbb{R}^{n}$ such that
  1. (i)
    $\dim(C_{\varepsilon})\leq\ell$;
  2. (ii)
    $C_{\varepsilon}\subset\mathcal{U}_{\varepsilon}(A)$;
  3. (iii)
    $\mathrm{dist}_{H}(F(C_{\varepsilon}),F(A))\leq L(F,r)\cdot\varepsilon$, where $L(F,r):=2+\mathrm{Lip}(F,B_{\mathbb{R}^{n}}(2r))$;
  4. (iv)
    for every $e=1,\ldots,n$ and every affine space $\mathbb{R}^{e}\simeq E\subset\mathbb{R}^{n}$, the number of connected components of $E\cap C_{\varepsilon}$ is bounded by
    (61) $b_{0}(E\cap C_{\varepsilon})\leq\beta_{\mathrm{dc}}^{e}.$
3.3. Variations and their behaviour under polynomial maps
We recall the definition of variations, introduced by Vitushkin [Vit55, Vit61] and developed in the semialgebraic context by Comte and Yomdin [YC04]. The variations encode the size of a set, through an integral-geometric approach. Given a polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ and a semialgebraic set $A\subset\mathbb{R}^{n}$, it is possible to estimate the variations of the set $p(A)$, in a quantitative way, see [YC04, Thm. 7.2]. However, these estimates are not quantitative with respect to the dimensional parameter $n$. The main result of this section is Theorem 18, where we prove a version of [YC04, Thm. 7.2], where the dependence on $n$ is explicit, thanks to the use of Theorem 13.
For $i\leq n$, we denote by $G_{i}(\mathbb{R}^{n})$ the Grassmannian of all $i$-dimensional linear subspaces of $\mathbb{R}^{n}$. We endow it with the standard probability measure $\gamma_{i,n}$ defined through the action of the orthogonal group of $\mathbb{R}^{n}$. For any $Z\in G_{i}(\mathbb{R}^{n})$ we the denote the orthogonal projection on $Z$ by
(62) $\pi_{Z}:\mathbb{R}^{n}\to Z.$
Definition 14 (Vitushkin variations).
Let $A\subset\mathbb{R}^{n}$ be a bounded semialgebraic set. We define the 00–th variation of $A$ as
(63) $V_{0}(A):=b_{0}(A),$
where $b_{0}(A)$ denotes the number of connected components of $A$. For $i=1,\ldots,n$ we define the $i$–th variation of $A$ as
(64) $V_{i}(A):=c(n,i)\int_{G_{i}(\mathbb{R}^{n})}\left(\int_{Z}b_{0}(A\cap\pi_{Z}^{-1}(x))\,\mathcal{H}^{i}(dx)\right)\,\gamma_{i,n}(dZ),$
where $\mathcal{H}^{i}$ denotes the $i$-dimensional Hausdorff measure.
(65) $c(n,i):=\frac{\Gamma(\frac{1}{2})\Gamma(\frac{n+1}{2})}{\Gamma(\frac{i+1}{2})\Gamma(\frac{n-i+1}{2})}.$
We note that the integrand in (64) is measurable, see [YC04, Sec. 3].
A list of properties of variations is given in [YC04, page 35]. Here we prove the following one.
Lemma 15.
If $A\subset\mathbb{R}^{n}$ is a bounded semialgebraic set of dimension $\ell$, then the $\ell$-variation coincides with the $\ell$-dimensional Hausdorff volume:
(66) $V_{\ell}(A)=\mathcal{H}^{\ell}(A).$
We also have $V_{i}(A)=0\text{ for all }i>\ell$.
Proof.
We recall a classical result in Integral Geometry, called Cauchy–Crofton formula [Fed47, Thm. 5.14]: if $A$ is a $i$-rectifiable and Borel subset of $\mathbb{R}^{n}$, for $i\leq n$, then it holds
(67) $\mathcal{H}^{i}(A)=c(n,i)\int_{G_{i}(\mathbb{R}^{n})}\left(\int_{Z}\#(A\cap\pi_{Z}^{-1}(x))\,\mathcal{H}^{i}(dx)\right)\,\gamma_{i,n}(dZ).$
If $A\subset\mathbb{R}^{n}$ is semialgebraic, the set $A$ can be written as a finite union of smooth, semialgebraic disjoint manifolds called strata (see Section 3.1.1). We can relabel the strata so that
(68) $A=\bigsqcup_{j=0}^{\ell}A_{j},$
where each $A_{j}$ is a finite union of smooth $j$-dimensional manifolds. For any fixed $Z\in G_{\ell}(\mathbb{R}^{N})$ and for almost every $x\in Z$ the set $\pi_{Z}^{-1}(x)$ does not intersect [^2] We note that in this step of the proof it is enough to use the Semialgebraic Sard’s theorem [BCR98, Thm. 9.6.2], which is logically independent from the result in the smooth category. $A_{j}$ for $j<\ell$. Hence, $V_{0}(A\cap\pi_{Z}^{-1}(x))=V_{0}(A_{\ell}\cap\pi_{Z}^{-1}(x))$ for almost every $x\in Z$. Furthermore, for almost every $x\in Z$ the intersection $A_{\ell}\cap\pi_{Z}^{-1}(x)$ is transverse, and $\dim A+\dim\pi_{Z}^{-1}(x)=n$, so that $A_{\ell}\cap\pi_{Z}^{-1}(x)$ has dimension zero. Thus, its cardinality coincides with its number of connected components. We conclude by (67). ∎
Let $M$ be a real $m\times m$ symmetric and positive semidefinite matrix. We denote by
(69) $\lambda_{1}(M)\geq\dots\geq\lambda_{m}(M)\geq 0,$
its eigenvalues, ordered in decreasing order.
Definition 16 (Singular values).
Let $H$ be a Hilbert space with $\dim H\geq m$, and let $L:H\to\mathbb{R}^{m}$ be a linear and continuous map. For any $k=1,\dots,m$ the $k$-th singular value of $L$ is
(70) $\sigma_{k}(L)=\lambda_{k}(LL^{\top})^{\frac{1}{2}}.$
In finite dimension, the following Weyl inequality holds for the singular values of the difference of two matrices (see [Tao12, Ex. 1.3.22 (iv)]): for $n\geq m$ and linear maps $L_{1},L_{2}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ we have
(71) $|\sigma_{k}(L_{1})-\sigma_{k}(L_{2})|\leq\|L_{1}-L_{2}\|_{\operatorname{op}},\qquad\forall\,k=1,\dots,m.$
The Weyl inequality generalizes to infinite-dimensional spaces. Since we are not able to find a reference, we provide below a self-contained statement and its proof in the form that we will need later.
Lemma 17.
If $L_{1},L_{2}\in\mathcal{L}(H,\mathbb{R}^{m})$, then $|\sigma_{k}(L_{1})-\sigma_{k}(L_{2})|\leq\|L_{1}-L_{2}\|_{\operatorname{op}}$ for all $k=1,\dots,m$.
Proof.
Let $L:H\to\mathbb{R}^{m}$ be linear and continuous. Let $W\subset H$ a finite dimensional subspace such that $(\ker L)^{\perp}\subseteq W$, and denote by $L_{W}:W\to\mathbb{R}^{m}$ the restriction of $L$ to $W$. More precisely, $L_{W}=L\circ i_{W}$ where $i_{W}:W\hookrightarrow H$ is the inclusion. We now prove that for all $k\leq m$
(72) $\sigma_{k}(L)=\sigma_{k}(L_{W}).$
Indeed, since the adjoint $i_{W}^{\top}$ coincides with the orthogonal projection $\pi_{W}:H\to W\subset H$, we have
(73) $\lambda_{k}(L_{W}\circ L_{W}^{\top})=\lambda_{k}(L\circ i_{W}\circ i_{W}^{\top}\circ L^{\top})=\lambda_{k}(L\circ\pi_{W}\circ L^{\top})=\lambda_{k}(L\circ L^{\top}).$
In the last equality, we used the fact that $L\circ\pi_{W}=L$, indeed $L(v)=L(\pi_{W}v+\pi_{W^{\perp}}v)=L(\pi_{W}v)$ since $W^{\perp}\subseteq\ker L$. Now given $L_{1},L_{2}\in\mathcal{L}(H,\mathbb{R}^{m})$ let $W:=(\ker L_{1})^{\perp}+(\ker L_{2})^{\perp}$, which is a finite dimensional linear subspace of $H$. By (72) and the finite dimensional Weyl inequality, we have
(74) $|\sigma_{k}(L_{1})-\sigma_{k}(L_{2})|=|\sigma_{k}((L_{1})_{W})-\sigma_{k}((L_{2})_{W})|\leq\|(L_{1})_{W}-(L_{2})_{W}\|_{\operatorname{op}}\leq\|L_{1}-L_{2}\|_{\operatorname{op}},$
concluding the proof. ∎
In the next result, given a semialgebraic set $A\subset\mathbb{R}^{n}$ and a polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$, we bound the $i$-th variation $V_{i}(p(A))$ in terms of the singular values of $p$ and the diagram of $A$. Its proof follows the blueprint of [YC04, Thm. 7.2]. The main novelty (which is key for our work) is that the dependence on the dimension $n$ is explicit.
Theorem 18.
For every $c,d,m\in\mathbb{N}$ there exists $\beta_{\mathrm{yc}}=\beta_{\mathrm{yc}}(c,d,m)>1$ with the following property. For $n\geq m$, let $A\subset\mathbb{R}^{n}$ be a closed and bounded semialgebraic set with $A\subset B_{\mathbb{R}^{n}}(r)$ and diagram $D(A)=(n,c,d)$. Let $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a polynomial map with components of degrees at most $d$. For all $i=1,\dots,m$ set
(75) $\overline{\sigma}_{i}:=\sup_{x\in A}\Big{(}{\sigma}_{1}(D_{x}p)\cdots{\sigma}_{i}(D_{x}p)\Big{)},\qquad\overline{\sigma}_{0}:=1.$
Then, for all $i=0,\dots,m$ it holds
(76) $V_{i}(p(A))\leq\mathrm{cst}(m,r)n^{m}\beta_{\mathrm{yc}}^{n}\overline{\sigma}_{i}$
where $\mathrm{cst}(m,r)>0$ is a constant that depends only on $m,r$.
Remark 1 (Origin of $\beta_{\mathrm{yc}}$).
An inspection of the proof shows that it holds $\beta_{\mathrm{yc}}=\beta_{\mathrm{tm}}^{2}\beta_{\mathrm{dc}}$ where $\beta_{\mathrm{tm}}=\beta_{\mathrm{tm}}(c(m+1),d)$ is the constant from the Thom-Milnor-type bound of Theorem 12, and $\beta_{\mathrm{dc}}=\max_{i\leq m}\beta_{\mathrm{dc}}(c,d,i)$ is the definable choice constant from Theorem 13.
Remark 2.
In the course of the proof, see (99), we obtain the following estimate, reminiscent of an area formula for variations, which is an improved version of [YC04, Thm. 7.2], with the explicit dependence of the constants on $n$: there exist a constant $\beta_{\mathrm{yc}}=\beta_{\mathrm{yc}}(c,d,m)$, such that
(77) $V_{i}(p(A))\leq c(m,i)\beta_{\mathrm{yc}}^{n}\overline{\sigma}_{i}V_{i}(A),\qquad i=1,\dots,m,$
where $c(m,i)$ is the constant in (65).
Proof.
By Definition 14, for $i=1,\dots,m$, we have
(78) $V_{i}(p(A))=c(m,i)\int_{G_{i}(\mathbb{R}^{m})}\left(\int_{Z}b_{0}(p(A)\cap\pi^{-1}_{Z}(x))\,\mathcal{H}^{i}(dx)\right)\gamma_{i,m}\left(dZ\right).$
For any $x\in Z$, the set $A\cap(\pi_{Z}\circ p)^{-1}(x)$ is semialgebraic with controlled diagram (more precisely, $D(A\cap(\pi_{Z}\circ p)^{-1}(x))=(n,c(1+m),d)$. Using Theorem 12 we obtain
(79) $b_{0}(p(A)\cap\pi^{-1}_{Z}(x))\leq b_{0}(A\cap(\pi_{Z}\circ p)^{-1}(x)))\leq\beta_{\mathrm{tm}}^{n},$
where $\beta_{\mathrm{tm}}=\beta_{\mathrm{tm}}(c(1+m),d)$ depends only on $c,d,m$. Note that, since by definition $V_{0}=b_{0}$, (79) yields the thesis of the theorem for the case $i=0$. We continue then assuming $i=1,\dots,m$. Denoting by $\mathbbm{1}_{S}$ the characteristic function of a set $S$, we get
(80) $V_{i}(p(A))$ $\leq c(m,i)\beta_{\mathrm{tm}}^{n}\int_{G_{i}(\mathbb{R}^{m})}\left(\int_{Z}\mathbbm{1}_{\pi_{Z}\circ p(A)}(x)\,\mathcal{H}^{i}(dx)\right)\,\gamma_{i,m}(dZ)$
(81) $=c(m,i)\beta_{\mathrm{tm}}^{n}\int_{G_{i}(\mathbb{R}^{m})}\mathcal{H}^{i}(\pi_{Z}\circ p(A))\,\gamma_{i,m}(dZ).$
We would like to estimate $\mathcal{H}^{i}(\pi_{Z}\circ p(A))$ applying area formula to the map
(82) $p_{Z}:=\pi_{Z}\circ p:\mathbb{R}^{n}\to\mathbb{R}^{i}.$
Exactly as in [YC04, Thm. 7.2], we cannot apply directly the area formula to $A$, since the dimension of $A$ can be larger than $i$. In [YC04, Thm. 7.2] this problem is solved using [YC04, Ex. 4.11]. However, the dependence on $n$ of the constants appearing there is not explicit. It could in be principle be made so, using the Tarski-Seidenberg theorem, but the dependence obtained in this way would be doubly exponential in $n$, which is not enough for our purposes. We overcome the obstacle using Theorem 13.
For sufficiently small $\varepsilon>0$ let $C_{\varepsilon}$ the semialgebraic set obtained by applying Theorem 13 to the set $A$ and the polynomial map $F=p_{Z}:\mathbb{R}^{n}\to\mathbb{R}^{i}$. We then proceed using the area formula to estimate $\mathcal{H}^{i}(\pi_{Z}\circ p(C_{\varepsilon}))$. (We will see at the end of the proof how $\mathcal{H}^{i}(\pi_{Z}\circ p(C_{\varepsilon}))\to\mathcal{H}^{i}(\pi_{Z}\circ p(A))$ as $\varepsilon\to 0$). By Item i of Theorem 13, $\dim(C_{\varepsilon})\leq i$. We can assume without loss of generality that $C_{\varepsilon}\subset\mathbb{R}^{n}$ is a smooth embedded submanifold with dimension equal to $i$, since the other strata give zero contribution to $\mathcal{H}^{i}$. Denoting with $\bar{p}_{Z}:C_{\varepsilon}\to\mathbb{R}^{i}$ the restriction, we obtain
(83) $\mathcal{H}^{i}(p_{Z}(C_{\varepsilon}))$ $\leq\int_{C_{\varepsilon}}|\det(D_{x}\bar{p}_{Z})|\mathcal{H}^{i}(dx)$ by the area formula
(84) $\leq\int_{C_{\varepsilon}}\sigma_{1}(D_{x}\bar{p}_{Z})\cdots\sigma_{i}(D_{x}\bar{p}_{Z})\mathcal{H}^{i}(dx).$ by definition of singular values
Note that, denoting with $j:C_{\varepsilon}\hookrightarrow\mathbb{R}^{n}$ the inclusion, we have
(85) $D_{x}\bar{p}_{Z}:T_{x}C_{\varepsilon}\xrightarrow{D_{x}j}\mathbb{R}^{n}\xrightarrow{D_{x}p}\mathbb{R}^{m}\xrightarrow{D_{p(x)}\pi_{Z}}\mathbb{R}^{i}.$
Thus, since $D_{p(x)}\pi_{Z}$ is an orthogonal projection and $D_{x}j$ is a linear inclusion, we have
(86) $\langle z,(D_{x}\bar{p}_{Z})(D_{x}\bar{p}_{Z})^{\top}z\rangle\leq\langle z,(D_{x}p)(D_{x}p)^{\top}z\rangle,\qquad\forall z\in Z=\mathbb{R}^{i}.$
It follows that $\sigma_{k}(D_{x}\bar{p}_{Z})\leq\sigma_{k}(D_{x}p)$ for all $k=1,\dots,i$. Continuing the above inequalities we obtain
(87) $\mathcal{H}^{i}(p_{Z}(C_{\varepsilon}))$ $\leq\int_{C_{\varepsilon}}\sigma_{1}(D_{x}p)\cdots\sigma_{i}(D_{x}p)\mathcal{H}^{i}(dx)$
(88) $\leq\sup_{x\in\mathcal{U}_{\varepsilon}(A)}\sigma_{1}(D_{x}p)\cdots\sigma_{i}(D_{x}p)\mathcal{H}^{i}(C_{\varepsilon})$ by Item ii of Theorem 13
(89) $\leq\left(\overline{\sigma}_{i}+\eta(\varepsilon)\right)\mathcal{H}^{i}(C_{\varepsilon}),$ by continuity of singular values
where $\overline{\sigma}_{i}$ is defined in (75), and $\eta(\varepsilon)\to 0$ as $\varepsilon\to 0$. Since $C_{\varepsilon}$ is a bounded semialgebraic set of dimension $i$, then $\mathcal{H}^{i}(C_{\varepsilon})=V_{i}(C_{\varepsilon})$ by Lemma 15. Hence
(90) $\mathcal{H}^{i}(C_{\varepsilon})$ $=c(n,i)\int_{G_{i}(\mathbb{R}^{n})}\left(\int_{Z}b_{0}(C_{\varepsilon}\cap\pi^{-1}_{Z}(x))\,\mathcal{H}^{i}(dx)\right)\gamma_{i,n}\left(dZ\right)$
(91) $\leq c(n,i)\beta_{\mathrm{dc}}^{n-i}\int_{G_{i}(\mathbb{R}^{n})}\left(\int_{Z}\mathbbm{1}_{\pi_{Z}(C_{\varepsilon})}(x)\,\mathcal{H}^{i}(dx)\right)\gamma_{i,n}\left(dZ\right)$ by Item iv of Theorem 13
(92) $\leq c(n,i)\beta_{\mathrm{dc}}^{n-i}\int_{G_{i}(\mathbb{R}^{n})}\left(\int_{Z}\mathbbm{1}_{\pi_{Z}(\mathcal{U}_{\varepsilon}(A))}(x)\,\mathcal{H}^{i}(dx)\right)\gamma_{i,n}\left(dZ\right)$ by Item ii of Theorem 13
(93) $\leq c(n,i)\beta_{\mathrm{dc}}^{n-i}\int_{G_{i}(\mathbb{R}^{n})}\left(\int_{Z}b_{0}(\mathcal{U}_{\varepsilon}(A)\cap\pi_{Z}^{-1}(x))\,\mathcal{H}^{i}(dx)\right)\gamma_{i,n}\left(dZ\right)$
(94) $=\beta_{\mathrm{dc}}^{n-i}V_{i}(\mathcal{U}_{\varepsilon}(A)),$
where $\beta_{\mathrm{dc}}=\beta_{\mathrm{dc}}(c,d,i)>1$ is the constant from Theorem 13, which (up to taking the maximum for $i\leq m$) depends only on $c,d,m$. Summing up, we have proved that for all $i=1,\dots,m$ it holds
(95) $\mathcal{H}^{i}(p_{Z}(C_{\varepsilon}))\leq(\overline{\sigma}_{i}+\eta(\varepsilon))\beta_{\mathrm{dc}}^{n}V_{i}(\mathcal{U}_{\varepsilon}(A)).$
In order to take the limit in (95), we recall two properties of semialgebraic sets. The first one is the continuity of the Lebesgue measure in the Hausdorff topology in the semialgebraic category (see [YC04, Thm. 5.10]): if $\{S_{\varepsilon}\}_{\varepsilon\geq 0}\subset\mathbb{R}^{i}$ is a one-parameter family of bounded semialgebraic sets then
(96) $\lim_{\varepsilon\to 0}\mathrm{dist}_{H}(S_{\varepsilon},S_{0})=0\qquad\Longrightarrow\qquad\lim_{\varepsilon\to 0}\mathcal{H}^{i}(S_{\varepsilon})=\mathcal{H}^{i}(S_{0}).$
The second property is the behaviour of variations for $\varepsilon$-neighbourhoods (see [YC04, Thm. 5.11]): if $A\subset\mathbb{R}^{n}$ is a bounded semialgebraic set, then for all $i=1,\dots,n$ it holds
(97) $\lim_{\varepsilon\to 0}V_{i}(\mathcal{U}_{\varepsilon}(A))\leq V_{i}(A).$
Therefore, using (96) (and recalling that by Item iii of Theorem 13 it holds $p_{Z}(C_{\varepsilon})\to p_{Z}(A)$ in the Hausdorff topology) and (97) we can pass to the limit in (95) and obtain
(98) $\mathcal{H}^{i}(p_{Z}(A))\leq\overline{\sigma}_{i}\beta_{\mathrm{dc}}^{n}V_{i}(A).$
Hence from (81) we get for all $i=1,\dots,m$
(99) $V_{i}(p(A))\leq c(m,i)\beta_{\mathrm{tm}}^{n}\beta_{\mathrm{dc}}^{n}\overline{\sigma}_{i}V_{i}(A).$
To conclude, we estimate $V_{i}(A)$. Similarly as in (81), and recalling that $A\subset B_{\mathbb{R}^{n}}(r)$ we have
(100) $V_{i}(A)\leq c(n,i)\beta_{\mathrm{tm}}^{n}\int_{G_{i}(\mathbb{R}^{n})}\mathcal{H}^{i}(\pi_{Z}(A))\gamma_{i,n}(dZ)\leq c(n,i)\beta_{\mathrm{tm}}^{n}\mathcal{H}^{i}(B_{\mathbb{R}^{i}}(r))=c(n,i)\beta_{\mathrm{tm}}^{n}\frac{\pi^{i/2}}{\Gamma\left(\frac{i}{2}+1\right)}r^{i}.$
Putting together (99) and (100), we obtain
(101) $V_{i}(p(A))\leq\left[c(m,i)c(n,i)\frac{\pi^{i/2}}{\Gamma\left(\frac{i}{2}+1\right)}\right](\beta_{\mathrm{tm}}^{2}\beta_{\mathrm{dc}})^{n}\overline{\sigma}_{i}r^{i}.$
Set $\beta_{\mathrm{yc}}=\beta_{\mathrm{tm}}^{2}\beta_{\mathrm{dc}}$. Using the form of the constants $c(m,i),c(n,i)$ in (65) and elementary estimates we obtain that for all $i=0,\dots,m$ there exists a constant $\mathrm{cst}(i)>0$ such that
(102) $V_{i}(p(A))\leq\mathrm{cst}(i)n^{i}\beta_{\mathrm{yc}}^{n}\overline{\sigma}_{i}r^{i}.$
Taking the maximum over $i=1,\dots,m$ we obtain the result with $\mathrm{cst}(m,r)=\displaystyle\max_{i=0,\dots,m}\mathrm{cst}(i)r^{i}$. ∎
3.4. Variations of almost-critical values of polynomial maps
In this section we recall the notion of almost-critical points $C^{\Lambda}(p)$ of a polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$. The set $p(C^{\Lambda}(p))$ is the set of its almost-critical values. The main result of this section is Theorem 23, which provides a version of [YC04, Cor. 7.4] where the dependence of the parameters with respect to $n$ is explicit. This result is based on Theorem 18 and on the specific semialgebraic structure of almost-critical points of polynomial maps, that we prove in Proposition 22.
Definition 19.
Let $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a $C^{1}$ map with $n\geq m$. Given $\Lambda=(\Lambda_{1},\dots,\Lambda_{m})\in\mathbb{R}^{m}_{+}$, the set of almost-critical points of $f$ is defined as
(103) $C^{\Lambda}(f):=\bigg{\{}x\in\mathbb{R}^{n}\bigg{|}\sigma_{i}(D_{x}f)\leq\Lambda_{i},\,\forall\,i=1,\dots,m\bigg{\}}.$
The following lemma is fundamental to study the semialgebraic structure of the set of almost-critical points of a polynomial map.
Lemma 20.
Let $A\subset\mathbb{R}^{s}$ be a semialgebraic set and let $f:A\to\mathbb{R}$ be a semialgebraic function. Then, for any $t\in\mathbb{R}$, the sublevel set $\{x\in A\mid f(x)\leq t\}$ admits a semialgebraic description whose diagram does not depend on $t$.
Proof.
The graph of $f$ is semialgebraic, hence we can write it as
(104) $\operatorname{graph}(f)=\bigcup_{i=1}^{a}\bigcap_{j=1}^{b_{i}}E_{ij},$
where $E_{ij}$ is of the form $\{p_{ij}<0\}$ or $\{p_{ij}\leq 0\}$ and $p_{ij}:\mathbb{R}^{s}\times\mathbb{R}\to\mathbb{R}$ is a polynomial. Denoting by $\pi_{1}:\mathbb{R}^{s}\times\mathbb{R}\to\mathbb{R}^{s}$ the projection, we have
(105) $\{x\in A\mid f(x)\leq t\}=\pi_{1}\Big{(}\operatorname{graph}(f)\cap\{(x,y)\in\mathbb{R}^{s}\times\mathbb{R}\mid y-t\leq 0\}\Big{)}.$
The set $\operatorname{graph}(f)\cap\{(x,y)\in\mathbb{R}^{s}\times\mathbb{R}\mid y-t\leq 0\}$ is semialgebraic, with a description given by
(106) $\operatorname{graph}(f)\cap\{(x,y)\mid y-t\leq 0\}=\bigcup_{i=1}^{a}\bigcap_{j=1}^{b_{i}+1}E_{i,j},$
where $E_{i,b_{i}+1}=\{(x,y)\in\mathbb{R}^{s}\times\mathbb{R}\mid y-t\leq 0\}$. In particular, this set admits a diagram that does not depend on $t$, since the degree of the polynomial $y-t$ is equal to $1$. As explained in Section 3.1.3, the projection of a semialgebraic set is semialgebraic and its diagram is controlled explicitly by the diagram of the original set only. In particular, the diagram of the projection (105) admits a description with a diagram $(s,I,J)$, that does not depend on $t$. ∎
Let $\operatorname{Sym}_{m}^{+}$ denote the set of positive semidefinite matrices of size $m$. This is clearly a semialgebraic subset of all $m\times m$ real matrices. The following lemma is elementary and we omit its proof.
Lemma 21.
The functions $\lambda_{i}:{\operatorname{Sym}}_{m}^{+}\to\mathbb{R}$ and are continuous and semialgebraic.
In the following proposition we provide an estimate on the diagram of a description of the (semialgebraic) set of almost-critical points of polynomial maps.
Proposition 22.
Let $d,m\in\mathbb{N}$. Then there are $c^{\prime}=c^{\prime}(d,m)$ and $d^{\prime}=d^{\prime}(d,m)$ such that for all $n\geq m$ and any polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with components of degrees at most $d$, and all $\Lambda=(\Lambda_{1},\dots,\Lambda_{m})\in\mathbb{R}^{m}_{+}$, the set $C^{\Lambda}(p)$ is closed, semialgebraic, with
(107) $D(C^{\Lambda}(p))=(n,c^{\prime},d^{\prime}).$
(In particular, the diagram does not depend on $\Lambda$ and it depends on $n$ only in the number of variables.)
Proof.
By Lemma 21, for every $k=1,\ldots,m$, the function $\lambda_{k}:{\operatorname{Sym}}_{m}^{+}\to\mathbb{R}$ is semialgebraic, hence from Lemma 20 applied to $A=\operatorname{Sym}_{m}^{+}$ and the semialgebraic functions $\lambda_{k}:A\to\mathbb{R}$, we deduce the following fact: for any $t\in\mathbb{R}$ we can write
(108) $\{M\in\operatorname{Sym}_{m}^{+}\mid\lambda_{k}(M)\leq t^{2}\}=\bigcup_{i=1}^{a}\bigcap_{j=1}^{b_{i}}\{M\in\mathbb{R}^{m\times m}\mid P_{k,i,j,t}(M)\leq 0\}$
for some polynomials $P_{k,i,j,t}:\mathbb{R}^{m\times m}\to\mathbb{R}$, $a,b_{1},\dots,b_{a}\in\mathbb{N}$, and with $a\max_{i}b_{i}$ and $\max_{i,j}\deg(P_{k,i,j,t})$ depending only on $m$ and on $k$ (and not on $t$). We can assume that $a$ does not depend on $k$, and all the $b_{i}$’s do not depend on $a,k$, simply by adding empty or trivial sets.
Now, for any $k=1,\dots,m$
(109) $\{N\in\mathbb{R}^{m\times n}\mid\sigma_{k}(N)\leq t\}=\{N\in\mathbb{R}^{m\times n}\mid\lambda_{k}(NN^{\top})\leq t^{2}\},$
therefore we can write
(110) $\{x\in\mathbb{R}^{n}\mid\sigma_{k}(D_{x}p)\leq t\}=\bigcup_{i=1}^{a}\bigcap_{j=1}^{b_{i}}\{x\in\mathbb{R}^{n}\mid P_{k,i,j,t}(D_{x}pD_{x}p^{\top})\leq 0\}.$
Observe that $N\to NN^{\top}$ is polynomial with components of degree $2$. Therefore, the map $q:\mathbb{R}^{n}\to\mathbb{R}^{m\times m},$ given by $q(x):=D_{x}pD_{x}p^{\top}$ is polynomial, with components of degree at most $2d$. Hence, the polynomial
(111) $P_{k,i,j,t}\circ q:\mathbb{R}^{n}\to\mathbb{R}$
has degree at most $2d\max_{i,j}\deg P_{k,i,j,t}$, and in particular its degree has a bound that depends on only $d,m,k$ (and not on $t$). The set of almost-critical points of $p$ can be described as
(112) $C^{\Lambda}(p)=\bigcap_{k=1}^{m}\{x\in\mathbb{R}^{n}\mid\sigma_{k}(D_{x}p)\leq\Lambda_{k}\}=\bigcap_{k=1}^{m}\bigcup_{i=1}^{a}\bigcup_{j=1}^{b_{i}}\{x\mid P_{k,i,j,\Lambda_{k}}(D_{x}pD_{x}p^{\top})\leq 0\}.$
It is clear that $C^{\Lambda}(p)$ is closed. Note that $d^{\prime}:=\max_{k,i,j}\deg(P_{k,i,j,\Lambda_{k}}\circ q)$ depends only on $d,m$. Using the distributivity of intersection, we can rewrite $C^{\Lambda}(p)$ in the form (47), where the number of unions and intersections depends only on $m$, and the maximum degree of the polynomials is $c^{\prime}=c^{\prime}(d,m)$, concluding the proof. ∎
Now we state and prove the main result of this section (Theorem G in the Introduction).
Theorem 23.
Let $n\geq m$, and $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a polynomial map with components of degree at most $d$. For $i=0,\dots,m$, $\Lambda=(\Lambda_{1},\dots,\Lambda_{m})\in\mathbb{R}^{m}_{+}$ and $r>0$, we have
(113) $V_{i}(p(C^{\Lambda}(p)\cap B_{\mathbb{R}^{n}}(r)))\leq\mathrm{cst}(m,r)n^{m}\beta_{0}^{n}\Lambda_{0}\cdots\Lambda_{i},$
where $\beta_{0}=\beta_{0}(d,m)>1$ depends only on $d$ and $m$, $\mathrm{cst}(m,r)$ depends only on $m,r$, and we set $\Lambda_{0}=1$.
Proof.
By Proposition 22 the set $A=C^{\Lambda}(p)\cap B_{\mathbb{R}^{n}}(r)$ is a closed and bounded semialgebraic set whose diagram is $D(A)=(n,c^{\prime},d^{\prime})$ where $c^{\prime}=c^{\prime}(d,m)$, $d^{\prime}=d^{\prime}(d,m)$ (which in particular do not depend on $n$, $\Lambda$). Furthermore, for the polynomial map $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ it holds, by construction
(114) $\overline{\sigma}_{i}:=\sup_{x\in A}\Big{(}{\sigma}_{1}(D_{x}p)\cdots{\sigma}_{i}(D_{x}p)\Big{)}\leq\Lambda_{0}\cdots\Lambda_{i}.$
We can then apply Theorem 18, yielding (113) with $\beta_{0}=\beta_{\mathrm{yc}}(c^{\prime},\max\{d,d^{\prime}\},m){\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}>1}$. ∎
4. Sard properties for infinite-dimensional maps
Let $H$ be a Hilbert space. We consider a map $f:U\to\mathbb{R}^{m}$ where $U$ is an open subset of $H$. If $f$ is $C^{1}$ we denote by $\operatorname{Crit}(f)$ the set of its critical points. For a fixed $\nu\in\mathbb{N}$, we also consider the set
(115) $\operatorname{Crit}_{\nu}(f)=\{x\in U\mid\mathrm{rank}(D_{x}f)\leq\nu\}.$
Given a relatively compact $S\subset\mathbb{R}^{m}$, for any $\varepsilon>0$ we denote with $M(\varepsilon,S)$ the $\varepsilon$-entropy of $S$, which is the minimum number of closed balls of radius $\varepsilon$ that we need to cover $S$.
Definition 24 (Entropy dimension).
The entropy dimension of $S$ is defined as
(116) $\dim_{e}(S)=\limsup_{\varepsilon\to 0^{+}}\frac{\ln M(\varepsilon,S)}{\ln(\frac{1}{\varepsilon})}.$
In [YC04, Ch. 2], it is proved that
(117) $\dim_{\mathcal{H}}(S)\leq\dim_{e}(S).$
The inequality can be strict, since there are sequences of real numbers with positive entropy dimension. We also note that the entropy dimension is stable under finite unions of sets, namely
(118) $\dim_{e}\left(\bigcup_{n=1}^{N}S_{n}\right)=\max_{n\leq N}\dim_{e}(S_{n}),$
while it is not even sub-additive with respect to countable unions. In fact, the entropy dimension of a point in $\mathbb{R}$ is equal to zero, while the entropy dimension of $\mathbb{Q}\subset\mathbb{R}$ is equal to one. (This is a major difference with respect to the Hausdorff dimension which is countably stable.)
For Banach spaces $X$, $Y$ and linear continuous maps $L\in\mathcal{L}(X,Y)$, we denote by
(119) $\|L\|_{\operatorname{op}}:=\sup_{\|x\|_{X}\leq 1}\|Lx\|_{Y}$
[…]
4.1. Sard-type theorem for well approximated maps
The purpose of this section is to prove the following theorem, which is our main Sard-type result (Theorem D in the Introduction).
Theorem 25.
Let $d,m\in\mathbb{N}$. There exists a constant $\beta_{0}=\beta_{0}(d,m)>1$ such that the following holds. Let $H$ be a Hilbert space, and let $f:H\to\mathbb{R}^{m}$ be a $\mathcal{C}^{1}$ map such that its differential $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous. Let $K\subset H$ be a bounded set with this approximation property: there exist a sequence $E_{n}\subset H$ of linear subspaces, $\dim(E_{n})=n$, and polynomial maps $f_{n}:E_{n}\to\mathbb{R}^{m}$ with uniformly bounded degree:
(120) $\sup_{n\in\mathbb{N}}\deg f_{n}\leq d<\infty,$
such that for some $q>1$, $c\geq 0$, and all large enough $n$ it holds
(121) $\sup_{x\in K}\bigg{(}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|+\|\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}\bigg{)}\leq cq^{-n}.$
Then for all $\nu\leq m-1$ it holds
(122) $\dim_{e}\bigg{(}f\big{(}\operatorname{Crit}_{\nu}(f)\cap K\big{)}\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q}.$
In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property on $K$:
(123) $\mu\bigg{(}f(\operatorname{Crit}(f)\cap K)\bigg{)}=0.$
Remark 3.
As it will be clear from the proof, the upper bound $cq^{-n}$ in the r.h.s. of (121) can be replaced by any $C_{n}\geq 0$, such that $\limsup_{n\to\infty}C_{n}^{1/n}=q^{-1}$, yielding the same results.
For the benefit of the reader we provide a general outline of the strategy of the proof. Our goal is to estimate the “size” of the set $f(\operatorname{Crit}_{\nu}(f)\cap K)$. More precisely, we bound the $\varepsilon$-entropy of $f(\operatorname{Crit}_{\nu}(f)\cap K)$ for any $\varepsilon>0$, and then pass to the limit $\varepsilon\to 0$.
We observe that the $\varepsilon$-entropy of $f(\operatorname{Crit}_{\nu}(f)\cap K)$ is controlled with the one of the set of almost-critical values $f(C^{\Lambda}(f)\cap K)$, see Definition 16. In turn, this is controlled by the $\varepsilon$-entropy of suitable almost-critical values of the approximating maps $f_{n}$, namely $f(C^{\Lambda_{\varepsilon}}(f_{n})\cap K)$, see (134).
To proceed, we use the relation between the $\varepsilon$-entropy of sets in $\mathbb{R}^{m}$ and their Vitushkin variations (Definition 14), provided by [YC04, Thm. 3.5].
Finally, we estimate the Vitushkin variations of $f_{n}(C^{\Lambda_{\varepsilon}}(f_{n})\cap K)$ applying Theorem G, exploiting the hypothesis that the maps $f_{n}$ are polynomial. This is the connection with quantitative semialgebraic geometry from Section 3. The crucial point of Theorem G is that the bound is quantitative with respect to parameters $\Lambda=\Lambda_{\varepsilon}$ and $n$. This allows to chose a large enough $n=n_{\varepsilon}$, see (148), to ensure convergence when passing to the limit $\varepsilon\to 0$.
Proof.
Step 1. We show that for any $\nu=0,\dots,m-1$ the set $\operatorname{Crit}_{\nu}(f)\cap K$ is contained in a suitable set of almost-critical points of $f$. By assumption, $K$ is weakly pre-compact. Since $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous, and by Lemma 17 the function $\sigma_{i}:\mathcal{L}(H,\mathbb{R}^{m})\to\mathbb{R}$ is continuous for every $i=1,\dots,m$. Therefore we have
(124) $\Sigma_{i}:=\sup_{K}\sigma_{i}(D_{x}f)<\infty.$
Hence, defining $\Lambda:=(\Sigma_{1},\dots,\Sigma_{\nu},0\dots,0)\in\mathbb{R}^{m}$, we get
(125) $\operatorname{Crit}_{\nu}(f)\cap K\subset C^{\Lambda}(f)\cap K.$
Step 2. We now relate critical points of $f$ to almost-critical points of the approximating polynomials $f_{n}$. Fix $\varepsilon>0$. Recall that, by Lemma 17, the singular values are $1$-Lipschitz with respect to the operator norm. Then for any sufficiently large $n\in\mathbb{N}$ such that $cq^{-n}\leq\varepsilon$ and for all $x\in K$ we have
(126) $|\sigma_{i}\left(D_{x}f|_{E_{n}}\right)-\sigma_{i}(D_{\pi_{E_{n}}(x)}f_{n})|\leq\|\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}\leq\varepsilon.$
Hence, for all $x\in K\cap C^{\Lambda}(f)$ we have the following estimate for all $i=1,\dots,m$
(127) $\sigma_{i}\left(D_{\pi_{E_{n}}(x)}f_{n}\right)\leq\varepsilon+\sup_{y\in C^{\Lambda}(f)\cap K}\sigma_{i}\left(D_{y}f|_{E_{n}}\right)\leq\varepsilon+\sup_{y\in C^{\Lambda}(f)\cap K}\sigma_{i}\left(D_{y}f\right)\leq\varepsilon+\Lambda_{i},$
where, in the second inequality, we used that singular values are monotone with respect to restriction to subspaces. Hence, defining
(128) $\Lambda^{\varepsilon}:=(\Sigma_{1}+\varepsilon,\dots,\Sigma_{\nu}+\varepsilon,\varepsilon,\dots,\varepsilon),$
by (127) we get
(129) $\pi_{E_{n}}\left(C^{\Lambda}(f)\cap K\right)\subset C^{\Lambda^{\varepsilon}}(f_{n})\cap\pi_{E_{n}}(K)\subset C^{\Lambda^{\varepsilon}}(f_{n})\cap B_{E_{n}}(r),$
where $r>0$ is such that $K\subseteq B_{H}(r)$, and thus $B_{E_{n}}(r)=B_{H}(r)\cap E_{n}$ denotes the ball in $E_{n}$ defined by the restriction of the Hilbert norm.
Step 3. Now we use the inclusion (129) to estimate the entropy dimension of $f(\operatorname{Crit}_{\nu}(f)\cap K)$ with the one of almost-critical values of $f_{n}$. By (121), for sufficiently large $n$, we have
(130) $\sup_{x\in K}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|\leq\varepsilon.$
Finally, we get
(131) $M(2\varepsilon,f(\operatorname{Crit}_{\nu}(f)\cap K)$ $\leq M(2\varepsilon,f(C^{\Lambda}(f)\cap K))$ by (125)
(132) $\leq M\left(2\varepsilon,\mathcal{U}_{\varepsilon}\left(f_{n}\circ\pi_{E_{n}}\left(C^{\Lambda}(f)\cap K\right)\right)\right)$ by (130)
(133) $\leq M\left(\varepsilon,f_{n}\circ\pi_{E_{n}}\left(C^{\Lambda}(f)\cap K\right)\right)$ by definition of $\mathcal{U}_{\varepsilon}$
(134) $\leq M(\varepsilon,f_{n}(C^{\Lambda^{\varepsilon}}(f_{n})\cap B_{E_{n}}(r))),$ by (129)
where we recall that $M(\varepsilon,A)$ is the minimal number of closed balls of $\mathbb{R}^{m}$ of radius $\varepsilon$ needed to cover $A\subset\mathbb{R}^{m}$, and $\mathcal{U}_{\varepsilon}(A)$ denotes the $\varepsilon$ neighbourhood of $A$. This concludes the proof of step 3.
The rest of the proof consists in estimating (134). We will use the results from Section 3.4, for this reason in the next step we discuss how to bring the problem to the Euclidean space $\mathbb{R}^{n}$.
Step 4. We fix a basis for $E_{n}$ which is orthonormal with respect to the inner product of $H$. This yields a linear isometry of Hilbert spaces $\ell_{n}:\mathbb{R}^{n}\to E_{n}$. In particular, for the transpose, it holds $\ell_{n}^{\top}=\ell_{n}^{-1}$. Consider then the polynomial map $\tilde{f}_{n}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ defined by $\tilde{f}_{n}:=f_{n}\circ\ell_{n}$, $n\geq m$. We have
(135) $\sigma_{i}(D_{v}\tilde{f}_{n})=\lambda_{i}(D_{\ell_{n}(v)}f_{n}\circ\ell_{n}\circ\ell_{n}^{\top}\circ D_{\ell_{n}(v)}f_{n}^{\top})^{1/2}=\sigma_{i}(D_{\ell_{n}(v)}f_{n}),\qquad\forall\,v\in\mathbb{R}^{n}.$
It follows that, for sufficiently large $n$,
(136) $\ell_{n}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n}))=C^{\Lambda^{\varepsilon}}(f_{n}).$
Hence, we have
(137) $f_{n}(C^{\Lambda^{\varepsilon}}(f_{n})\cap B_{E_{n}}(r))=f_{n}(\ell_{n}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n}))\cap B_{E_{n}}(r))=\tilde{f}_{n}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n})\cap B_{\mathbb{R}^{n}}(r)).$
By [YC04, Thm. 3.5] we have
(138) $M(\varepsilon,\tilde{f}_{n}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n})\cap B_{\mathbb{R}^{n}}(r)))\leq C(m)\sum_{i=0}^{m}\frac{1}{\varepsilon^{i}}V_{i}\bigg{(}\tilde{f}_{n}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n})\cap B_{\mathbb{R}^{n}}(r))\bigg{)},$
where $C(m)>0$ is a constant that depends only on $m$. We now apply the results from Section 3.4 in order to estimate the right-hand side of the last inequality.
Step 5. Take $n$ sufficiently large so that it also satisfies $n\geq m$. Then, by Theorem G, for every $i=0,\dots,m$ we have
(139) $V_{i}\bigg{(}\tilde{f}_{n}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n})\cap B_{\mathbb{R}^{n}}(r))\bigg{)}\leq\mathrm{cst}(m,r)n^{m}\beta_{0}^{n}\Lambda_{0}^{\varepsilon}\cdots\Lambda_{i}^{\varepsilon},$
where we recall $\beta_{0}=\beta_{0}(d,m)$ and $\Lambda_{0}^{\varepsilon}=1$ by definition. Now, we have for all $i=1,\dots,m$
(140) $\Lambda^{\varepsilon}_{1}\cdots\Lambda^{\varepsilon}_{i}$ $=(\Lambda_{1}+\varepsilon)\cdots(\Lambda_{i}+\varepsilon)$
(141) $=\sum_{h=0}^{i}\varepsilon^{i-h}\sum_{0=j_{0}<j_{1}<\dots<j_{h}\leq i}\Lambda_{j_{0}}\cdots\Lambda_{j_{h}}$
(142) $\leq\sum_{h=0}^{i}\varepsilon^{i-h}\sum_{0=j_{0}<j_{1}<\cdots<j_{h}\leq i}\Lambda_{0}\cdots\Lambda_{h}$
(143) $=\binom{i}{h}\sum_{h=0}^{i}\varepsilon^{i-h}\Lambda_{0}\cdots\Lambda_{h}$
(144) $\leq i!\sum_{h=0}^{i}\varepsilon^{i-h}\Lambda_{0}\cdots\Lambda_{h}.$
Therefore, by (138), (139) and (144), for large $n$ we have
(145) $M(\varepsilon,\tilde{f}_{n}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n})\cap B_{\mathbb{R}^{n}}(r)))$ $\leq\mathrm{cst}(m,r)C(m)n^{m}\beta_{0}^{n}\sum_{i=0}^{m}\frac{i!}{\varepsilon^{i}}\sum_{h=0}^{i}\varepsilon^{i-h}\Lambda_{0}\cdots\Lambda_{h}$
(146) $\leq\widetilde{\mathrm{cst}}(m,r)n^{m}\beta_{0}^{n}\sum_{h=0}^{\nu}\frac{\Lambda_{0}\cdots\Lambda_{h}}{\varepsilon^{h}},$
where we note that, by definition, $\Lambda_{h}=0$ for all $h>\nu$, and $\widetilde{\mathrm{cst}}(m,r)$ denotes a constant that depends only on $m,r$. From this we deduce that
(147) $\ln M(\varepsilon,\tilde{f}_{n}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n})\cap B_{\mathbb{R}^{n}}(r)))\leq\ln(\widetilde{\mathrm{cst}}(m,r))+n\ln\beta_{0}+m\ln n+\ln\left(\sum_{h=0}^{\nu}\frac{\Lambda_{0}\cdots\Lambda_{h}}{\varepsilon^{h}}\right),$
for all $\varepsilon>0$ and $n\in\mathbb{N}$ such that $cq^{-n}\leq\varepsilon$ and $n\geq m$.
Step 6. We now show how to use (147) in order to get the estimate on $\dim_{e}(f(\operatorname{Crit}_{\nu}(f)\cap K))$. For $\varepsilon>0$ we choose $n=n_{\varepsilon}$ where
(148) $n_{\varepsilon}:=\left\lceil\log_{q}\frac{c}{\varepsilon}\right\rceil,$
in such a way that (134), (137) and (147) hold when $\varepsilon\to 0$. By (134) and (137) we have
(149) $\dim_{e}f(\operatorname{Crit}_{\nu}(f)\cap K)\leq\limsup_{\varepsilon\to 0}\frac{\ln M(\varepsilon,\tilde{f}_{n_{\varepsilon}}(C^{\Lambda^{\varepsilon}}(\tilde{f}_{n_{\varepsilon}})\cap B_{\mathbb{R}^{{n_{\varepsilon}}}}(r)))}{\ln\frac{1}{2\varepsilon}}.$
Hence now we estimate the right-hand side through (147). As $\varepsilon$ goes to zero we have
(150) $\limsup_{\varepsilon\to 0}\frac{n_{\varepsilon}\ln\beta_{0}}{\ln\frac{1}{2\varepsilon}}=\lim_{\varepsilon\to 0}\frac{\left\lceil\log_{q}\frac{c}{\varepsilon}\right\rceil\ln\beta_{0}}{\ln\frac{1}{2\varepsilon}}=\frac{\ln\beta_{0}}{\ln q}.$
We also have, taking into account (148), that
(151) $\lim_{\varepsilon\to 0}\frac{m\ln n_{\varepsilon}}{\ln(\frac{1}{2\varepsilon})}=0.$
Furthermore
(152) $\lim_{\varepsilon\to 0}\frac{1}{\ln(\frac{1}{2\varepsilon})}\ln\left(\sum_{h=0}^{\nu}\frac{\Lambda_{0}\cdots\Lambda_{h}}{\varepsilon^{h}}\right)=\nu.$
Therefore,
(153) $\dim_{e}f(\operatorname{Crit}_{\nu}(f)\cap K)\leq\nu+\frac{\ln\beta_{0}}{\ln q},$
concluding the proof. ∎
4.2. Sard-type theorems and Kolmogorov $n$-width
In this section, we recall the definition of Kolmogorov $n$–width, and we show its role in Sard-type theorems.
Definition 26 (Kolmogorov $n$–width).
Let $X$ be a normed space and $S\subset X$ be a subset. For $n\in\mathbb{N}$, the Kolmogorov $n$–width of $S$ in $X$ is
(154) $\Omega_{n}(S,X)=\inf_{\dim E=n}\sup_{u\in S}\inf_{y\in E}\|u-y\|,$
where the infimum is taken over all $n$-dimensional linear subspaces $E$ of $X$.
The asymptotics of $n$–width measures in a quantitative way the compactness of a set, in fact the following holds (see [Pin85, Prop. 1.2]).
Proposition 27.
$S$ is compact if and only if $S$ is bounded and $\lim_{n\to\infty}\Omega_{n}(S)=0$.
For various properties of the $n$–width we refer the reader to [Pin85]; we recall here those that we will need in the sequel, see [Pin85, Thm. 1.1].
Theorem 28.
Let $X$ be a normed space and $S\subset X$. Then for all $n\in\mathbb{N}$:
  1. (i)
    $\Omega_{n}(S,X)=\Omega_{n}(\bar{S},X)$, where $\bar{S}$ is the closure of $S$;
  2. (ii)
    For every $\alpha\in\mathbb{R}$
    (155) $\Omega_{n}(\alpha S,X)=|\alpha|\Omega_{n}(S,X);$
  3. (iii)
    Let $\mathrm{co}(S)$ be the convex hull of $S$. Then
    (156) $\Omega_{n}(\mathrm{co}(S),X)=\Omega_{n}(S,X);$
  4. (iv)
    Let $\mathrm{b}(S)=\{\alpha x\mid x\in S,\,|\alpha|\leq 1\}$ be the balanced hull of $S$. Then
    (157) $\Omega_{n}(\mathrm{b}(S),X)=\Omega_{n}(S,X).$
Remark 29.
As a consequence, in all our results, up to enlarging $S$, one can assume without loss of generality that $S$ is convex and centrally symmetric, without changing its $n$–width.
We establish now a connection between $n$–width of compact sets and locally Lipschitz functions.
Lemma 30.
Let $H$ be a Hilbert space, $(Y,\|\cdot\|_{Y})$ a Banach space, and let $f:H\to Y$ be locally Lipschitz. Let $K\subset H$ be a compact subset. Then there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ there exist a linear subspace $E_{n}\subset H$ of dimension $n$ such that
(158) $\sup_{x\in K}\|f(x)-f(\pi_{E_{n}}(x))\|_{Y}\leq c(f,K)\Omega_{n}(K,H).$
Proof.
Let $\{B_{H}(x_{i},r/2)\}_{i\in I}$ a finite cover of $K$ by balls, with centers $x_{i}\in K$ and radii $r/2>0$. We can assume that $f$ is $L$-Lipschitz on each $B_{H}(x_{i},r)$, for some $L>0$. Since $K$ is compact $\Omega_{n}(K,H)\to 0$ as $n\to\infty$, hence there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ we have $\Omega_{n}(K,H)<r/4$. By definition of $n$–width, for all $n\geq n_{0}$ there exists a $n$-dimensional subspace $E_{n}\subset H$ such that
(159) $\sup_{x\in K}\|x-\pi_{E_{n}}(x)\|<2\Omega_{n}(K,H)<\frac{r}{2}.$
Let $x\in K$. Then by construction $x\in B_{H}(x_{i},r/2)$ for some $i\in I$. By (159) we have $\pi_{E_{n}}(x)\in B_{H}(x_{i},r)$. Since $f$ is $L$-Lipschitz on every $B_{H}(x_{i},r)$ we have
(160) $\sup_{x\in K}\|f(x)-f(\pi_{E_{n}}(x))\|_{Y}\leq L\sup_{x\in K}\|x-\pi_{E_{n}}(x)\|\leq 2L\Omega_{n}(K,H),$
where we used again (159). This proves the result, with $c(f,K)=2L$. ∎
We can now prove the following Sard-type theorem, corresponding to Theorem A in the Introduction.
Theorem 31.
Let $H$ be a Hilbert space and $K\subset H$ be a compact set such that, for some $q>1$, it holds
(161) $\limsup_{n\to\infty}\Omega_{n}(K,H)^{1/n}\leq q^{-1}.$
Let $d,m\in\mathbb{N}$. There exists $\beta_{0}=\beta_{0}(d,m)>1$ such that for all $f\in\mathscr{P}_{d}^{m}(H)$ and $\nu\leq m-1$ we have
(162) $\dim_{e}\bigg{(}f\big{(}\operatorname{Crit}_{\nu}(f)\cap K\big{)}\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q}.$
In particular, if $q>\beta_{0}$, then the Sard property holds on $K$:
(163) $\mu\bigg{(}f\big{(}\operatorname{Crit}(f)\cap K\big{)}\bigg{)}=0.$
Proof.
Let $q_{\varepsilon}\in(0,q)$. By assumption, $\Omega_{n}(K,H)\leq q^{-n}_{\varepsilon}$ for sufficiently large $n$. We prove that $f$ satisfies the hypothesis of Theorem D. Since $f\in\mathscr{P}^{m}_{d}(H)$ (see Definition 1), the map
(164) $(f,Df):H\to\mathbb{R}^{m}\times\mathcal{L}(H,\mathbb{R}^{m})$
is locally Lipschitz. The codomain $Y=\mathbb{R}^{m}\times\mathcal{L}(H,\mathbb{R}^{m})$ is a Banach space equipped with the norm $\|(v,A)\|_{Y}=\|v\|+\|A\|_{\operatorname{op}}$ for $v\in\mathbb{R}^{m}$ and $A\in\mathcal{L}(H,\mathbb{R}^{m})$. Hence we can apply Lemma 30 to get $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ there exists a $n$-dimensional linear subspace $E_{n}\subset H$ such that
(165) $\sup_{x\in K}\|f(x)-f(\pi_{E_{n}}(x))\|+\|D_{x}f-D_{{\pi_{E_{n}(x)}}}f\|_{\operatorname{op}}\leq c(f,K)q^{-n}_{\varepsilon},$
where $\pi_{E_{n}}:H\to E_{n}$ denotes the orthogonal projection. Let $f_{n}:E_{n}\to\mathbb{R}^{m}$ defined by the restriction $f_{n}:=f|_{E_{n}}$, which are polynomial maps with degree uniformly bounded above by $d$. It holds
(166) $\|(D_{x}f)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}=\|(D_{x}f)|_{E_{n}}-(D_{\pi_{E_{n}}(x)}f)|_{E_{n}}\|_{\operatorname{op}}\leq\|D_{x}f-D_{\pi_{E_{n}}(x)}f\|_{\operatorname{op}}.$
As a consequence of the above two estimates, assumption (18) of Theorem D holds, yielding
(167) $\dim_{e}\bigg{(}f\big{(}\operatorname{Crit}_{\nu}(f)\cap K\big{)}\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q_{\varepsilon}}.$
Letting $q_{\varepsilon}\uparrow q$ we obtain the thesis. ∎
4.3. Counterexamples to the Sard theorem in infinite dimension
In this section we provide examples of maps $f\in\mathscr{P}^{m}_{d}(H)$ as in Theorem A for which there exists a compact set $K\subset H$ such that the set $f(\operatorname{Crit}(f)\cap K)$ has not measure zero, however it holds $\Omega_{n}(K,H)\leq cq^{-n}$. This shows the necessity of the quantitative assumption $q>\beta_{0}(m,d)$. As a consequence, we get lower bounds on the semialgebraic constant $\beta_{0}$.
We start with the following example, which is a minor modification of Kupka’s one [Kup65].
Example 32 (Kupka revisited).
Let $q>1$. Let $f:\ell^{2}\to\mathbb{R}$ defined by
(168) $f(x)=\sum_{k=1}^{\infty}\frac{1}{2^{k}}\phi(q^{k-1}x_{k}),$
where $\phi$ is the Kupka polynomial as in [YC04, Sec. 10.2.3]. Namely $\phi$ has degree $3$, with $\phi(0)=\phi^{\prime}(0)=\phi(1)-1=\phi^{\prime}(1)=0$. If we assume $q^{3}/2<1$ then $f$ is $C^{\infty}$. All critical points have the form
(169) $\operatorname{Crit}(f)=\left\{\left(x_{1},\frac{x_{2}}{q},\dots,\frac{x_{k}}{q^{k-1}},\dots\right)\,\middle|\,x_{i}\in\{0,1\}\right\},$
and the set of critical values is $[0,1]$. Furthermore, given $y=\left(x_{1},\frac{x_{2}}{q},\dots,\frac{x_{k}}{q^{k-1}},\dots\right)\in\operatorname{Crit}(f)$, we consider its projection $y_{n}$ on $\mathrm{span}\{e_{1},\dots,e_{n}\}$, that is
(170) $y_{n}=\left(x_{1},\frac{x_{2}}{q},\dots,\frac{x_{n}}{q^{n-1}},0,0,0,\dots\right).$
We have
(171) $\left\lVert y-y_{n}\right\rVert_{\ell^{2}}^{2}\leq\sum_{k=n+1}^{\infty}\frac{1}{q^{2(k-1)}}\leq\int_{n}^{\infty}\frac{1}{q^{2t}}=\frac{q^{-2n}}{2\ln q}.$
Hence,
(172) $\Omega_{n}(\operatorname{Crit}(f),\ell^{2})\leq\sup_{y\in\operatorname{Crit}(f)}\left\lVert y-y_{n}\right\rVert_{\ell^{2}}\leq\frac{q^{-n}}{\sqrt{2\ln q}}\qquad\Longrightarrow\qquad\limsup_{n\to\infty}\Omega_{n}(\operatorname{Crit}(f),\ell^{2})\leq q^{-1}.$
Since $f$ and $\operatorname{Crit}(f)$ satisfy all the hypothesis of Theorem A (provided that $q^{3}/2<1$), we get a consequence that
(173) $\beta_{0}(1,3)\geq 2^{1/3}.$
We now get the lower bound on $\beta_{0}(1,d)$ for $d\geq 3$ by improving on the above construction.
Example 33 (Kupka revisited, degree $d\geq 3$).
We need the following classical fact, which follows from [Kuh69]: for any $d\geq 2$ there exists a polynomial $\psi:\mathbb{R}\to\mathbb{R}$ of degree $d$ such that
(174) $\psi(\operatorname{Crit}(\psi))=\{0,\dots,d-2\}.$
(See also [Kri84, Thm. 1] which gives a version of this result taking into account also possible multiplicities.) We fix $d\geq 3$ and we consider a polynomial $\psi$ of degree $d$ such that (174) holds. For $q>1$ we consider the map $f_{d}:\ell^{2}\to\mathbb{R}$
(175) $f_{d}(x):=\sum_{k=1}^{\infty}\frac{1}{(d-1)^{k}}\psi(q^{k-1}x_{k}).$
If $q\in(1,(d-1)^{1/d})$, then the map $f_{d}$ is $C^{\infty}$. The set of critical points of $f_{d}$ is
(176) $\operatorname{Crit}(f_{d})=\left\{\left(y_{1},\frac{y_{2}}{q},\dots,\frac{y_{k}}{q^{k-1}},\dots\right)\in\ell^{2}\,\middle|\,\psi^{\prime}(y_{k})=0,\,\forall\,k\in\mathbb{N}\right\}.$
Hence we have
(177) $f_{d}(\operatorname{Crit}(f_{d}))=\left\{\sum_{k=1}^{\infty}\frac{1}{(d-1)^{k}}c_{k}\,\middle|\,c_{k}\in\{0,\dots,d-2\}\right\}=[0,1].$
Let $\zeta=\max\{|z|\mid\psi^{\prime}(z)=0\}$. Exactly as in Example 32 we get for every $n\in\mathbb{N}$
(178) $\Omega_{n}(\operatorname{Crit}(f_{d}),\ell^{2})\leq\frac{\zeta}{\sqrt{2\ln q}}q^{-n}\qquad\Longrightarrow\qquad\limsup_{n\to\infty}\Omega_{n}(\operatorname{Crit}(f_{d}),\ell^{2})\leq q^{-1}.$
For any $q\in(1,(d-1)^{1/d})$ the function $f_{d}$ and the set $K=\operatorname{Crit}(f_{d})$ satisfy the hypothesis of Theorem 31, but $f_{d}(\operatorname{Crit}(f_{d}))=[0,1]$. It follows that
(179) $\beta_{0}(1,d)\geq(d-1)^{1/d},\qquad\forall\,d\geq 3.$
In the next example we extend the construction to arbitrary dimension of the codomain.
Example 34 (Kupka revisited, $d\geq 3$, arbitrary codomain).
Let $f_{d}$ be as in Example 33, and consider $g_{d}:\mathbb{R}^{m-1}\times\ell^{2}\to\mathbb{R}^{m}$ defined as
(180) $g_{d}(x,y):=(x,f_{d}(y)).$
We have
(181) $\operatorname{Crit}(g_{d})=\mathbb{R}^{m-1}\times\operatorname{Crit}(f_{d}).$
Taking $E_{n+m-1}=\mathrm{span}\{(e_{i},0),\,(0,e_{j})\mid i=1,\dots,m-1,\,j=1,\dots,n\}\subset\mathbb{R}^{m-1}\times\ell^{2}$, for all $n\in\mathbb{N}$, we can estimate the $n$–width of $\operatorname{Crit}(g_{d})$ as
(182) $\Omega_{n+m-1}(\operatorname{Crit}(g_{d}),\mathbb{R}^{m-1}\times\ell^{2})\leq\frac{\zeta}{\sqrt{2\ln q}}q^{-n}\qquad\Longrightarrow\qquad\limsup_{n\to\infty}\Omega_{n}(\operatorname{Crit}(g_{d}),\mathbb{R}^{m-1}\times\ell^{2})\leq q^{-1}.$
We observe, provided that $q\in(1,(d-1)^{1/d})$, the function $g_{d}$ satisfies all the hypothesis of Theorem 31, but the set $g_{d}(\operatorname{Crit}(g_{d}))=\mathbb{R}^{m-1}\times[0,1]$ has not measure zero. Hence
(183) $\beta_{0}(d,m)\geq(d-1)^{1/d},\qquad\forall\,m\in\mathbb{N},\,d\geq 3.$
We collect the content of Examples 32, 33 and 34 in a unified statement (see Theorem B).
Theorem 35.
Let $d,m\in\mathbb{N}$, with $d\geq 3$, and let $1<q<(d-1)^{1/d}$. There exist a Hilbert space $H$, and $f\in\mathscr{P}_{d}^{m}(H)$ such that $K=\operatorname{Crit}(f)\cap B_{H}(r)$ is compact for all $r>0$, with
(184) $\limsup_{n\to\infty}\Omega_{n}(K,H)^{1/n}\leq q^{-1},$
and $f:H\to\mathbb{R}^{m}$ does not verify the Sard property, namely $\mu(f(\operatorname{Crit}(f)\cap K))>0$. Therefore, the semialgebraic constant $\beta_{0}(d,m)$ of Theorem A satisfies
(185) $\beta_{0}(d,m)\geq(d-1)^{1/d},\qquad\forall\,m\in\mathbb{N},\,d\geq 3.$
In all the examples above we have taken $K$ to be set of all critical points of a given function in a ball. We now show a different construction, where the set $K$ is strictly contained in the set of critical points, more precisely is the set of points where the differential has rank zero.
Example 36 (Rank zero counterexample to Sard).
We consider $f_{d}:\ell^{2}\to\mathbb{R}$ as in Example 33, and $m\in\mathbb{N}$. We define $h:(\ell^{2})^{m}\to\mathbb{R}^{m}$ as
(186) $h(x^{1},\dots,x^{m}):=(f_{d}(x^{1}),\dots,f_{d}(x^{m})).$
For any $x=(x^{1},\dots,x^{m})\in(\ell^{2})^{m}$ we have
(187) $\operatorname{im}(D_{x}h)=\operatorname{im}(D_{x^{1}}f_{d})\times\dots\times\operatorname{im}(D_{x^{m}}f_{d}).$
Hence $x\in\operatorname{Crit}_{\nu}(h)$ (the set of points where the rank of the differential is $\leq\nu$) if and only if at least $m-\nu$ components of $x$ are critical points of $f_{d}$. In particular, we consider the compact set $K$ of all critical points of $h$ of rank zero, namely
(188) $K:=\operatorname{Crit}_{0}(h)=\operatorname{Crit}(f_{d})^{m}.$
We have $K\subsetneq\operatorname{Crit}(h)$, and
(189) $h(\operatorname{Crit}(h)\cap K)=h(\operatorname{Crit}_{0}(h))=[0,1]^{m}.$
As in the previous examples, $K$ is compact with exponential $n$–width. By taking finite-dimensional spaces $E_{nm}=\mathrm{span}\{\underbrace{(0,\dots,e_{i},\dots,0)}_{j\text{-th component}}\in(\ell^{2})^{m}\mid i=1,\dots,n,\,j=1,\dots,m\}$, for $n\in\mathbb{N}$ we have
(190) $\Omega_{nm}(K,(\ell^{2})^{m})\leq m\Omega_{n}(\operatorname{Crit}(f_{d}),\ell^{2})\leq\frac{m\zeta}{\sqrt{2\ln q}}q^{-n}\qquad\Longrightarrow\qquad\limsup_{n\to\infty}\Omega_{n}(K,(\ell^{2})^{m})\leq q^{-1/m},$
where $\zeta$ is the same constant appearing in Example 33.
4.4. Sard on linear subspaces
Next, we prove Theorem C, of which we recall the statement.
Theorem 37.
Let $H$ be a Hilbert space and $K\subset H$ be a compact set such that, for some $q>1$, it holds
(191) $\limsup_{n\to\infty}\Omega_{n}(K,H)^{1/n}\leq q^{-1}.$
Let $d,m\in\mathbb{N}$, let $\beta_{0}=\beta_{0}(d,m)>1$ be the same constant of Theorem A. Consider the (possibly non-closed) linear subspace
(192) $V:=\mathrm{span}(K).$
Then for all $f\in\mathscr{P}_{d}^{m}(H)$ and $\nu\leq m-1$ the restriction $f|_{V}:V\to\mathbb{R}^{m}$ satisfies
(193) $\dim_{\mathcal{H}}\bigg{(}f\big{(}\mathrm{Crit}_{\nu}(f|_{V})\big{)}\bigg{)}\leq\nu+\frac{\log\beta_{0}}{\log q}.$
In particular, if $q>\beta_{0}$, then the restriction $f|_{V}:V\to\mathbb{R}^{m}$ satisfies the Sard property
(194) $\mu\bigg{(}f\big{(}\mathrm{Crit}(f|_{V})\big{)}\bigg{)}=0.$
In all previous results, $f\big{(}\mathrm{Crit}(f|_{V})\big{)}$ can be replaced with the smaller set $f(\mathrm{Crit}(f)\cap V)$.
Proof.
Let $\tilde{H}$ denote the closure of $V$ in $H$. Observe first that
(195) $\Omega_{n}(K,H)=\Omega_{n}(K,\tilde{H}).$
The inequality $\Omega_{n}(K,H)\leq\Omega_{n}(K,\tilde{H})$ is clear since every $n$-dimensional subspace of $\tilde{H}$ is also a subspace of $H$. For the other inequality, since $\tilde{H}$ is closed in $H$, we can write $H=\tilde{H}\oplus\tilde{H}^{\perp}$ and denote by $\pi:H\to\tilde{H}$ the orthogonal projection. Then, using that $K\subset V\subset\tilde{H}$, we obtain
(196) $\Omega_{n}(K,H)$ $=\inf_{\begin{subarray}{c}Y\subset H\\ \dim Y=n\end{subarray}}\sup_{u\in K}\inf_{y\in Y}\|u-y\|\geq\inf_{\begin{subarray}{c}Y\subset H\\ \dim Y=n\end{subarray}}\sup_{u\in K}\inf_{y\in Y}\|u-\pi(y)\|$
(197) $=\inf_{\begin{subarray}{c}Y\subset H\\ \dim Y=n\end{subarray}}\sup_{u\in K}\inf_{z\in\pi(Y)}\|u-z\|$
(198) $=\inf_{\begin{subarray}{c}Z\subset\tilde{H}\\ \dim Z\leq n\end{subarray}}\sup_{u\in K}\inf_{z\in Z}\|u-z\|$
(199) $=\inf_{\begin{subarray}{c}Z\subset\tilde{H}\\ \dim Z=n\end{subarray}}\sup_{u\in K}\inf_{z\in Z}\|u-z\|=\Omega_{n}(K,\tilde{H}).$
We can apply Theorem A to the map $\tilde{f}\in\mathscr{P}_{d}^{m}(\tilde{H})$, defined by $\tilde{f}:=f|_{\tilde{H}}$, with $f\in\mathscr{P}_{d}^{m}(H)$, and the compact set $K\subset\tilde{H}$ obtaining for all $\nu\leq m-1$
(200) $\dim_{e}\bigg{(}\tilde{f}\big{(}\operatorname{Crit}_{\nu}(\tilde{f})\cap K\big{)}\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q}.$
Observe now that if $u$ is a critical point for $f|_{V}$, then $u$ is also critical for $\tilde{f}=f|_{\tilde{H}}$. In fact, $D_{u}(f|_{V})=(D_{u}f)|_{V}$ for all $u\in V$. Thus we have $\lambda\in(\mathbb{R}^{m})^{*}\setminus\{0\}$ such that $0\equiv\lambda\circ D_{u}(f|_{V})=(\lambda\circ D_{u}\tilde{f})|_{V}$ and, since $V$ is dense in $\tilde{H}$, it follows that $\lambda\circ D_{u}\tilde{f}\equiv 0$. This means that
(201) $\mathrm{Crit}_{\nu}(f|_{V})\cap K\subseteq\mathrm{Crit}_{\nu}(\tilde{f})\cap K.$
In particular, since on $K\subset V$ we have $\tilde{f}=f|_{V}$ by definition, this implies that
(202) $\dim_{e}\bigg{(}f\big{(}\operatorname{Crit}_{\nu}(f|_{V})\cap K\big{)}\bigg{)}\leq\dim_{e}\bigg{(}\tilde{f}\big{(}\operatorname{Crit}_{\nu}(\tilde{f})\cap K\big{)}\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q}.$
Let $K^{\prime}$ be the balanced and convex hull of $K$. It holds:
(203) $V=\mathrm{span}(K)=\mathrm{span}(K^{\prime})=\bigcup_{j\in\mathbb{N}}jK^{\prime}.$
By Theorem 28 the $n$–width of the compact, convex, centrally symmetric sets $jK^{\prime}$ satisfy
(204) $\Omega_{n}(jK^{\prime},H)=j\Omega_{n}(K^{\prime},H)=j\Omega_{n}(K,H).$
Therefore we can repeat the above argument, obtaining (202) for $jK^{\prime}$ in place of $K$, for any $j\in\mathbb{N}$. Furthermore, we can replace the entropy dimension with the (smaller) Hausdorff dimension, which has the advantage of being countably stable. Taking the supremum over $j\in\mathbb{N}$ we obtain (193).
To conclude, observe that any critical point of $f|_{V}$ is also a critical point of $f$ that is also in $V$, thus $f(\mathrm{Crit}(f|_{V}))\subseteq f(\mathrm{Crit}(f)\cap V)$, which yields the last claim in the statement of the theorem. ∎
4.5. A class of maps with the global Sard property
Using Theorem D, we study a class of maps from a Hilbert space to $\mathbb{R}^{m}$ for which the Sard property holds true. The case $m=1$ is not new, and it was proved in [Yom88], see also [YC04]. We need the following preliminary facts.
Lemma 38 (Markov inequality).
Let $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a polynomial map with components of degree at most $d$. Then for every $r>0$ it holds
(205) $\sup_{x\in B_{\mathbb{R}^{n}}(r)}\|D_{x}p\|_{\operatorname{op}}\leq\frac{\sqrt{m}d^{2}}{r}\sup_{x\in B_{\mathbb{R}^{n}}(r)}\|p(x)\|.$
Proof.
From [Kel28, Thm. VI] we get the thesis for $r=1$. Then, we conclude considering the rescaled polynomial $x\to p(rx)$. ∎
Lemma 39 (Estimates on balls at different radii).
Let $p:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a polynomial map with components of degree at most $d$. Then for all $r>0$ it holds
(206) $\sup_{x\in B_{\mathbb{R}^{n}}(r)}\|p(x)\|\leq\alpha(d,m)n^{d}(1+r)^{d}\sup_{x\in B_{\mathbb{R}^{n}}(1)}\|p(x)\|,$
where $\alpha(d,m)=m^{d/2}d^{2d}(d+1)$ is a constant depending only on $d$ and $m$.
Proof.
We can write the polynomial map as
(207) $p(x)=\sum_{|\alpha|\leq d}c_{\alpha}x^{\alpha},$
where $c_{\alpha}\in\mathbb{R}^{m}$, and the sum is over the multi-indices $\alpha\in\mathbb{N}^{n}$ with $|\alpha|=\sum_{i}\alpha_{i}\leq d$. Since $|\alpha|!c_{\alpha}=\partial_{x^{\alpha}}p|_{x=0}$, we can iterate Markov inequality (for $r=1$, see Lemma 38) and obtain
(208) $\|c_{\alpha}\|\leq\frac{1}{|\alpha|!}m^{|\alpha|/2}d^{2|\alpha|}\sup_{x\in B_{\mathbb{R}^{n}}(1)}\|p(x)\|\leq m^{|\alpha|/2}d^{2|\alpha|}\sup_{x\in B_{\mathbb{R}^{n}}(1)}\|p(x)\|.$
Therefore, for $r>1$, we have
(209) $\sup_{x\in B_{\mathbb{R}^{n}}(r)}\|p(x)\|$ $=\sup_{x\in B_{\mathbb{R}^{n}}(1)}\|p(rx)\|$
(210) $=\sup_{x\in B_{\mathbb{R}^{n}}(1)}\left\|\sum_{i=1}^{d}r^{i}\left(\sum_{|\alpha|=i}c_{\alpha}x^{\alpha}\right)\right\|$
(211) $\leq(1+r)^{d}\left(\sum_{|\alpha|\leq d}m^{|\alpha|/2}d^{2|\alpha|}\right)\sup_{x\in B_{\mathbb{R}^{n}}(1)}\|p(x)\|$
(212) $\leq(1+r)^{d}m^{d/2}d^{2d}(d+1)n^{d}\sup_{x\in B_{\mathbb{R}^{n}}(1)}\|p(x)\|,$
where we used the fact that the number of multi-indices $\alpha\in\mathbb{N}^{n}$ with $|\alpha|\leq d$ is equal to $(1+n+n^{2}+\dots+n^{d})\leq(d+1)n^{d}$
For the rest of this section, we assume that the Hilbert space $H$ is separable, and we fix a Hilbert basis $\{e_{i}\}_{i\in\mathbb{N}}\subset H$. For $k\in\mathbb{N}$, we set $E_{k}:=\mathrm{span}\{e_{1},\ldots,e_{k}\}$ and we denote by $\pi_{k}:H\to E_{k}$ the corresponding orthogonal projection, that is if $x=\sum_{i}x_{i}e_{i}$, then $\pi_{k}(x)=(x_{1},\dots,x_{k})$. We now define a special class of maps in this setting.
Proposition 40 (Construction of special maps).
Let $H$ be a separable Hilbert space. For all $k\in\mathbb{N}$, let $p_{k}:E_{k}\to\mathbb{R}^{m}$ be polynomial maps with $\sup_{k\in\mathbb{N}}\deg p_{k}\leq d$ for some $d\in\mathbb{N}$, and such that
(213) $\sum_{k=1}^{\infty}\sup_{x\in B_{E_{k}}(r)}\|p_{k}(x)\|<\infty,\qquad\forall\,r>0.$
Then the map $f:H\to\mathbb{R}^{m}$ given by
(214) $f(x):=\sum_{k=1}^{\infty}p_{k}(x_{1},\dots,x_{k}),\qquad\forall\,x\in H,$
is well–defined, $f\in\mathscr{P}_{d}^{m}(H)$ (see Definition 1) and its differential is given by
(215) $D_{x}f=\sum_{k=1}^{\infty}D_{x}(p_{k}\circ\pi_{k}).$
Proof.
For a Banach spaces $(X,\|\cdot\|_{X})$, $(Y,\|\cdot\|_{Y})$, and a closed set $U\subseteq X$, we denote by $\mathcal{B}(U,Y)$ the space of all continuous and bounded functions $f:U\to Y$. It is a Banach space when endowed with the norm $\sup_{x\in U}\left\lVert f(x)\right\rVert_{Y}$.
We first note that the assumption (213) implies that the sequence of maps $f_{n}:H\to\mathbb{R}^{m}$ given by $f_{n}:=\sum_{k=1}^{n}p_{k}\circ\pi_{k}$ is Cauchy in $\mathcal{B}(B_{H}(r),\mathbb{R}^{m})$ for all $r>0$. In particular the right hand side of (214) converges locally uniformly. For any $n\in N$ we have
(216) $D_{x}f_{n}=\sum_{k=1}^{n}D_{x}(p_{k}\circ\pi_{k}).$
Furthermore, $x\mapsto D_{x}f_{n}\in\mathcal{B}(B_{H}(r);\mathcal{L}(H,\mathbb{R}^{m}))$ for all $n\in\mathbb{N}$, $r>0$. We now prove that it is a Cauchy sequence. Indeed, using Lemma 38, for any $n_{2}\geq n_{1}$
(217) $\sup_{x\in B_{H}(r)}\left\lVert D_{x}f_{n_{2}}-D_{x}f_{n_{1}}\right\rVert_{\operatorname{op}}\leq\frac{\sqrt{m}d^{2}}{r}\sum_{k>n_{1}}\sup_{x\in B_{E_{k}}(r)}\|p_{k}(x)\|\to 0,$
as $n_{2},n_{1}\to\infty$. Therefore, $\{Df_{n}\}_{n\in\mathbb{N}}$ has a limit in $\mathcal{B}(B_{H}(r),\mathcal{L}(H,\mathbb{R}^{m}))$, for all $r>0$. It follows that $f$ is $\mathcal{C}^{1}$ and (215) holds, see e.g. [Lan93, Thm. 9.1].
Thanks to (215), we prove that $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is weakly continuous. Let us consider a sequence $\{x^{j}\}_{j\in\mathbb{N}}$ weakly convergent to $x\in B_{H}(r)$; we prove that $D_{x^{j}}f\to D_{x}f$ as $j\to\infty$. Indeed, given $\varepsilon>0$ we fix $N_{\varepsilon}\in\mathbb{N}$ such that $\sum_{k>N_{\varepsilon}}\sup_{B_{E_{k}}(r)}\|p_{k}\|\leq\varepsilon$ and, by Lemma 38, we get
(218) $\left\lVert D_{x^{j}}f-D_{x}f\right\rVert_{\operatorname{op}}$ $\leq\sum_{k=1}^{\infty}\left\lVert D_{x^{j}}(p_{k}\circ\pi_{k})-D_{x}(p_{k}\circ\pi_{k})\right\rVert_{\operatorname{op}}$
(219) $=\sum_{k=1}^{\infty}\left\lVert(D_{\pi_{k}(x^{j})}p_{k})\circ\pi_{k}-(D_{\pi_{k}(x)}p_{k})\circ\pi_{k}\right\rVert_{\operatorname{op}}$
(220) $=\sum_{k=1}^{\infty}\left\lVert D_{\pi_{k}(x^{j})}p_{k}-D_{\pi_{k}(x)}p_{k}\right\rVert_{\operatorname{op}}$
(221) $\leq\sum_{k=1}^{N_{\varepsilon}}\left\lVert D_{\pi_{k}(x^{j})}p_{k}-D_{\pi_{k}(x)}p_{k}\right\rVert_{\operatorname{op}}+\sum_{k>N_{\varepsilon}}\left\lVert D_{\pi_{k}(x^{j})}p_{k}\right\rVert_{\operatorname{op}}+\sum_{k>N_{\varepsilon}}\left\lVert D_{\pi_{k}(x)}p_{k}\right\rVert_{\operatorname{op}}$
(222) $\leq\sum_{k=1}^{N_{\varepsilon}}\left\lVert D_{\pi_{k}(x^{j})}p_{k}-D_{\pi_{k}(x)}p_{k}\right\rVert_{\operatorname{op}}+\frac{\sqrt{m}d^{2}}{r}2\varepsilon,$
where with some abuse of notation we denoted by the same symbol $\|\cdot\|_{\operatorname{op}}$ either the norm in $\mathcal{L}(H,\mathbb{R}^{m})$ or the one for the finite-dimensional subspaces $\mathcal{L}(E_{k},\mathbb{R}^{m})$. Since $\pi_{k}(x^{j})\to\pi_{k}(x)$ as $j\to\infty$ for all $k\in\mathbb{N}$, we conclude by the continuity of $Dp_{k}$ on $E_{k}$.
We prove now that $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is locally Lipschitz. We will prove that the sequence $\{Df_{n}\}_{n\in\mathbb{N}}$ defined by (216) is uniformly Lipschitz on balls. For $k\in\mathbb{N}$, fix $z,w\in E_{k}$. Then it holds
(223) $D_{z}p_{k}-D_{w}p_{k}=D_{z}p_{k}-D_{z}\tilde{p}_{k}=D_{z}(p_{k}-\tilde{p}_{k}),$
where $\tilde{p}_{k}:E_{k}\to\mathbb{R}^{m}$, defined by $\tilde{p}_{k}(\cdot):=p_{k}(\cdot-z+w)$, is polynomial depending on the fixed choices of $z,w$, with degree bounded by $d$. Thus for all fixed $z,w\in B_{E_{k}}(r)$, $r>0$, it holds
(224) $\|D_{z}p_{k}-D_{w}p_{k}\|_{\operatorname{op}}$ $\leq\sup_{a\in B_{E_{k}}(r)}\|D_{a}(p_{k}-\tilde{p}_{k})\|_{\operatorname{op}}$ by (223)
(225) $\leq\frac{\sqrt{m}d^{2}}{r}\sup_{a\in B_{E_{k}}(r)}\|p_{k}(a)-\tilde{p}_{k}(a)\|$ by Lemma 38
(226) $\leq\frac{\sqrt{m}d^{2}}{r}\sup_{a\in B_{E_{k}}(r)}\|p_{k}(a)-p_{k}(a-z+w)\|$ by definition of $\tilde{p}_{k}$
(227) $\leq\left(\frac{\sqrt{m}d^{2}}{r}\sup_{a\in B_{E_{k}}(3r)}\|D_{a}p_{k}\|_{\operatorname{op}}\right)\|z-w\|$ $p_{k}$ is Lipschitz
(228) $\leq\left(\frac{md^{4}}{r^{2}}\sup_{a\in B_{E_{k}}(3r)}\|p_{k}(a)\|\right)\|z-w\|.$ by Lemma 38
Note that if $x\in B_{H}(r)$ then $\pi_{k}(x)\in B_{E_{k}}(r)$. Thus, for all $x,y\in B_{H}(r)$ it holds
(229) $\|D_{x}f_{n}-D_{y}f_{n}\|_{\operatorname{op}}$ $\leq\sum_{k=1}^{n}\|D_{x}(p_{k}\circ\pi_{k})-D_{y}(p_{k}\circ\pi_{k})\|_{\operatorname{op}}$ by (216)
(230) $\leq\sum_{k=1}^{n}\|D_{\pi_{k}(x)}p_{k}-D_{\pi_{k}(y)}p_{k}\|_{\operatorname{op}}$ by Leibniz rule
(231) $\leq\frac{md^{4}}{r^{2}}\sum_{k=1}^{n}\sup_{a\in B_{E_{k}}(3r)}\|p_{k}(a)\|\|\pi_{k}(x)-\pi_{k}(y)\|$ by (228)
(232) $\leq\frac{md^{4}}{r^{2}}\|x-y\|\sum_{k=1}^{\infty}\sup_{a\in B_{E_{k}}(3r)}\|p_{k}(a)\|$ $\pi_{k}$ is $1$-Lipschitz
(233) $\leq c(r)\|x-y\|,$
where $c(r)>0$ does not depend on $n$, and in the last inequality we used (213) for the convergence of the series. Therefore the sequence $\{Df_{n}\}_{n\in\mathbb{N}}$, is uniformly Lipschitz on every ball $B_{H}(r)$. Thus its limit $Df:H\to\mathcal{L}(H,\mathbb{R}^{m})$ is Lipschitz on any ball $B_{H}(r)$, and a fortiori locally Lipschitz.
Finally, we show that for every finite dimensional space $E\subset H$, the map $f|_{E}$ is a polynomial map with components of degree bounded by $d$. Let $n=\mathrm{dim}(E)$ and fix a linear isometry $L:\mathbb{R}^{n}\to E$. The statement is equivalent to show that $f\circ L$ is a polynomial map with components of degree bounded by $d$. Denoting by $(t_{1},\ldots,t_{n})$ the standard coordinates in $\mathbb{R}^{n}$, we have
(234) $f(L(t_{1},\ldots,t_{n}))$ $=\sum_{k=1}^{\infty}p_{k}(x_{1}(t_{1},\ldots,t_{n}),\dots,x_{k}(t_{1},\ldots,t_{n}))$
(235) $=\sum_{k=1}^{\infty}\tilde{p}_{k}(t_{1},\ldots,t_{n}),$
where each $\tilde{p}_{k}$ is a polynomial in $(t_{1},\ldots,t_{n})$ of degree at most $d$ (since the $x_{i}$ depend linearly on the $t_{j}$). Condition (213) guarantees that, for each monomial $t_{1}^{\alpha_{1}}\cdots{t_{n}}^{\alpha_{n}}$, its coefficient in the expansion $\sum_{k\leq s}\tilde{p}_{k}$ converges as $s\to\infty$. Note that there are only a finite number of such coefficients, determined by the upper bound $d$ on the degree and the dimension $n$. We conclude by noting that the uniform limit of polynomials with degree bounded by $d$ is a polynomial with degree bounded by $d$. ∎
With the following we exhibit a special class of maps, obtained via the construction of Proposition 40, for which the Sard property holds true globally. This is the content of Theorem E in the Introduction, of which we recall the statement here.
Theorem 41.
Let $H$ be a separable Hilbert space. For all $k\in\mathbb{N}$, let $p_{k}:E_{k}\to\mathbb{R}^{m}$ be polynomial maps with $\sup_{k\in\mathbb{N}}\deg p_{k}\leq d$ for some $d\in\mathbb{N}$, and such that
(236) $\sup_{x\in B_{E_{k}}(1)}\|p_{k}(x)\|\leq q^{-k},\qquad\forall\,k\in\mathbb{N},$
for some $q>1$. Then the map $f:H\to\mathbb{R}^{m}$ defined by
(237) $f(x):=\sum_{k=1}^{\infty}p_{k}(x_{1},\dots,x_{k}),\qquad\forall\,x\in H,$
is well–defined, $f\in\mathscr{P}_{d}^{m}(H)$ (see Definition 1), and for all $\nu\leq m-1$ and $r>0$ it holds
(238) $\dim_{e}\bigg{(}f\left(\operatorname{Crit}_{\nu}(f)\cap B_{H}(r)\right)\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q},$
where $\beta_{0}=\beta_{0}(d,m)>1$ is the same constant given by Theorem D. In particular, if $q>\beta_{0}$, then $f$ satisfies the Sard property globally on $H$:
(239) $\mu\bigg{(}f\big{(}\operatorname{Crit}(f)\big{)}\bigg{)}=0.$
Proof.
We show first that the map $f$ is well–defined. The upper bound (236), together with Lemma 39 imply that for all $r>0$
(240) $\sup_{x\in B_{E_{k}}(r)}\|p_{k}(x)\|\leq\alpha(d,m)(1+r)^{d}\frac{k^{d}}{q^{k}},\qquad\forall\,k\in\mathbb{N},$
with $q>1$. In turn, this implies that
(241) $\sum_{k=1}^{\infty}\sup_{x\in B_{E_{k}}(r)}\|p_{k}(x)\|<\infty,\qquad\forall\,r>0.$
We can then apply Proposition 40 obtaining that $f\in\mathscr{P}_{d}^{m}(H)$.
Fix $r>0$. We now prove that $f$ satisfies the hypothesis of Theorem D with $K=B_{H}(r)$ and polynomial approximating maps $f_{n}:E_{n}\to\mathbb{R}^{m}$
(242) $f_{n}(x):=\sum_{k=1}^{n}p_{k}(x_{1},\dots,x_{k}),\qquad\forall\,x\in H.$
Note that, by our assumptions, $\sup_{n\in\mathbb{N}}\deg f_{n}\leq d$. It remains to check the validity of (18). Firstly, we have for all $n\in\mathbb{N}$
(243) $\sup_{K}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|\leq\alpha(d,m)(1+r)^{d}\sum_{k=n+1}^{\infty}\frac{k^{d}}{q^{k}},$
where we used (240).
We now estimate the derivatives. We have
(244) $\sup_{x\in K}\left\lVert\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\right\rVert_{\operatorname{op}}$ $\leq\sup_{x\in K}\left\lVert D_{x}f-D_{x}(f_{n}\circ\pi_{E_{n}})\right\rVert_{\operatorname{op}}$
(245) $\leq\sup_{x\in K}\sum_{k=n+1}^{\infty}\left\lVert D_{x}(p_{k}\circ\pi_{k})\right\rVert_{\operatorname{op}}$ by (215)
(246) $=\sup_{x\in K}\sum_{k=n+1}^{\infty}\left\lVert(D_{\pi_{k}(x)}p_{k})\circ\pi_{k}\right\rVert_{\operatorname{op}}$
(247) $\leq\sum_{k=n+1}^{\infty}\sup_{x\in B_{E_{k}}(r)}\left\lVert D_{x}p_{k}\right\rVert_{\operatorname{op}}.$
Note that, in the last line, the supremum is taken on a finite-dimensional Euclidean ball. To proceed, by Lemma 38 and (240) we have
(248) $\sup_{x\in B_{E_{k}}(r)}\left\lVert D_{x}p_{k}\right\rVert_{\operatorname{op}}\leq\frac{d^{2}\sqrt{m}}{r}\alpha(d,m)(1+r)^{d}\frac{k^{d}}{q^{k}}.$
Therefore, continuing the previous estimate, we obtain
(249) $\sup_{x\in K}\left\lVert\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\right\rVert_{\operatorname{op}}$ $\leq\frac{d^{2}\sqrt{m}}{r}\alpha(d,m)(1+r)^{d}\sum_{k=n+1}^{\infty}\frac{k^{d}}{q^{k}}.$
We have therefore proved that for all $n\in\mathbb{N}$, and $r>0$ it holds
(250) $\sup_{x\in K}\bigg{(}\|f(x)-f_{n}\circ\pi_{E_{n}}(x)\|+\|\left(D_{x}f\right)|_{E_{n}}-D_{\pi_{E_{n}}(x)}f_{n}\|_{\operatorname{op}}\bigg{)}\leq\left(1+\frac{d^{2}\sqrt{m}}{r}\right)\alpha(d,m)(1+r)^{d}\sum_{k=n+1}^{\infty}\frac{k^{d}}{q^{k}}.$
To conclude, note that by elementary estimates it holds
(251) $\left(1+\frac{d^{2}\sqrt{m}}{r}\right)\alpha(d,m)(1+r)^{d}\sum_{k=n+1}^{\infty}\frac{k^{d}}{q^{k}}\leq cq^{-n},$
for sufficiently large $n$, where $c\geq 1$ is a constant depending only on the fixed parameters $r,d,m$. Therefore, the main assumption (18) of Theorem D is satisfied, yielding
(252) $\dim_{e}\bigg{(}f\big{(}\operatorname{Crit}_{\nu}(f)\cap B_{H}(r)\big{)}\bigg{)}\leq\nu+\frac{\ln\beta_{0}}{\ln q},\qquad\forall\,\nu={\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}1}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0},\dots,m-1.$
In particular, if $q>\beta_{0}$ then $f$ satisfies the Sard property on $B_{H}(r)$, namely
(253) $\mu\bigg{(}f\big{(}\operatorname{Crit}(f)\cap B_{H}(r)\big{)}\bigg{)}=0.$
Since $r>0$ is arbitrary, we obtain that $\mu(f(\operatorname{Crit}(f)))=0$. ∎
4.6. Sard threshold on compacts
We can now prove Theorem F, of which we recall the statement.
Theorem 42.
For every $d,m\in\mathbb{N}$, there exists $\omega_{0}(d,m)\in(0,1]$ such that
  1. (i)
    for every $f\in\mathscr{P}_{d}^{m}(H)$ and for every compact set $K\subset H$ with $\omega(K,H)<\omega_{0}(d,m)$,
    (254) $\mu\bigg{(}f\left(\mathrm{crit}(f)\cap K\right)\bigg{)}=0;$
  2. (ii)
    for every $\omega>\omega_{0}(d,m)$, with $\omega\in(0,1]$, there exist $f\in\mathscr{P}_{d}^{m}(H)$ and a compact set $K\subset H$ with $\omega(K,H)=\omega$ and such that
    (255) $\mu\bigg{(}f\left(\mathrm{crit}(f)\cap K\right)\bigg{)}>0.$
Proof.
By definition, for any compact set $K$, it holds $\omega(K,H)\in[0,1]$, see (25). We define $\omega_{0}(d,m)$ as the supremum of the set of $\omega\in[0,1]$ such that for every $f\in\mathscr{P}_{d}^{m}(H)$ and for every compact set $K\subset H$ with $\omega(K,H)=\omega$ it holds:
(256) $\mu\bigg{(}f\left(\mathrm{Crit}(f)\cap K\right)\bigg{)}=0.$
By Theorem A this set is non-empty and $\omega_{0}(d,m)\geq\beta_{0}(d,m)^{-1}$. Hence $\omega_{0}(d,m)\in(0,1]$. Both items of Theorem 42 follow. Note that by Theorem B (more specifically, the construction of Example 34), provided that $d\geq 3$, one must have $\omega_{0}(d,m)\leq(d-1)^{-1/d}<1$, so that Item ii is non-vacuous. ∎
5. Applications to the Endpoint maps of Carnot groups
We apply in this last section the results from the previous sections to the study of the Sard property for Endpoint maps for Carnot groups.
Definition 43 (Carnot groups).
An $m$-dimensional Carnot group of step $s\in\mathbb{N}$ is a connected and simply connected Lie group $(\mathbb{G},\cdot)$ of dimension $m$, whose Lie algebra of left-invariant vector fields $\mathfrak{g}$ admits a stratification of step $s$, that is
(257) $\mathfrak{g}=\mathfrak{g}_{1}\oplus\dots\oplus\mathfrak{g}_{s},$
where $\mathfrak{g}_{i}\neq\{0\}$, $\mathfrak{g}_{i+1}=[\mathfrak{g}_{1},\mathfrak{g}_{i}]$ for all $i=1,\dots,s-1$ and $[\mathfrak{g}_{1},\mathfrak{g}_{s}]=\{0\}$. We set $k_{i}=\dim\mathfrak{g}_{i}$, $a_{i}=k_{1}+\dots+k_{i}$, and $k=k_{1}$, which is called the rank of the Carnot group.
We identify elements of $\mathfrak{g}$ with vectors of $T_{e}\mathbb{G}$. Since $\mathfrak{g}$ is nilpotent and $\mathbb{G}$ is simply connected, the group exponential map $\exp_{\mathbb{G}}:\mathfrak{g}\to\mathbb{G}$ is a smooth diffeomorphism.
The stratification (257) allows the definition of a family of automorphisms $\delta_{\lambda}:\mathbb{G}\to\mathbb{G}$, for $\lambda>0$, called dilations such that
(258) $\delta_{\lambda}(\exp_{\mathbb{G}}(v))=\exp_{\mathbb{G}}(\lambda^{i}v),\qquad\forall\,v\in\mathfrak{g}_{i},\quad i=1,\dots,s.$
Fix an adapted basis of $\mathfrak{g}$, namely a set of vectors $v_{1},\dots,v_{m}\in\mathfrak{g}$ such that for all $i=1,\dots,s$ the list of vectors $v_{k_{i-1}+1},\dots,v_{k_{i}}$ is a basis of $\mathfrak{g}_{i}$, with the convention $k_{0}=0$. For any $j=1,\dots,m$ the weight of $v_{j}$ is the unique $w_{j}\in\{1,\dots,s\}$ such that $v_{j}\in\mathfrak{g}_{w_{j}}$. We define a diffeomorphism $\varphi_{\mathbb{G}}:\mathbb{R}^{m}\to\mathbb{G}$, by
(259) $\varphi_{\mathbb{G}}(x_{1},\dots,x_{m}):=\exp_{\mathbb{G}}\left(\sum_{i=1}^{m}x_{i}v_{i}\right).$
The map $\varphi_{\mathbb{G}}$ defines global coordinates on $\mathbb{G}$, called exponential coordinates (of first kind). For any $x\in\mathbb{G}$ we denote by $\tau_{x}:\mathbb{G}\to\mathbb{G}$ the left translation on $\mathbb{G}$, that is $\tau_{x}(y)=x\cdot y$. We denote by $X_{1},\dots,X_{m}$ the left-invariant vector fields on $\mathbb{G}$ corresponding to the adapted basis. They are given by $X_{j}(x)=D_{e}\tau_{x}(v_{j})$ for any $x\in\mathbb{G}$, where $D_{e}\tau_{x}$ denotes the differential of $\tau_{x}$ at the unit element.
We recall the definition of (homogeneous) polynomials on Carnot groups. We say that $P:\mathbb{G}\to\mathbb{R}$ is a polynomial if for some (and then any) set of exponential coordinates $\varphi_{\mathbb{G}}$, the map $P\circ\varphi_{\mathbb{G}}:\mathbb{R}^{m}\to\mathbb{R}$ is a polynomial. Any polynomial $P:\mathbb{G}\to\mathbb{R}$ can be written as
(260) $P(\varphi_{\mathbb{G}}(x_{1},\dots,x_{m}))=\sum_{\alpha\in\mathbb{N}^{m}}c_{\alpha}x_{1}^{\alpha_{1}}\cdots x_{m}^{\alpha_{m}},\qquad\forall\,x\in\mathbb{R}^{m},$
for some $c_{\alpha}\in\mathbb{R}$, which are non-zero only for a finite set of multi-indices. For $\alpha\in\mathbb{N}^{m}$, we denote by
(261) $|\alpha|_{\mathbb{G}}:=\sum_{i=1}^{m}w_{i}\alpha_{i},$
the weighted degree of the multi-index $\alpha$. The weighted degree of a polynomial $P$ is
(262) $\deg_{\mathbb{G}}(P):=\max\left\{|\alpha|_{\mathbb{G}}\,\middle|\,c_{\alpha}\neq 0\right\}.$
A polynomial $P$ is homogeneous of weighted degree $w$ if $P(\delta_{\lambda}x)=\lambda^{w}P(x)$ for any $x\in\mathbb{G}$ and $\lambda>0$.
Remark 4.
In other words, $P$ is homogeneous of weighted degree $w$ if and only if
(263) $P(\varphi_{\mathbb{G}}(x_{1},\dots,x_{m}))=\sum_{\begin{subarray}{c}\alpha\in\mathbb{N}^{m}\\ |\alpha|_{\mathbb{G}}=w\end{subarray}}c_{\alpha}x_{1}^{\alpha_{1}}\cdots x_{m}^{\alpha_{m}}.$
In particular any such polynomial depends only on the variables $x_{i}$ with $w_{i}\leq w$, namely $x_{1},\dots,x_{k_{w}}$.
It is easy to see that the concepts we just introduced do not depend on the choice of exponential coordinates. For this reason, in the following, we fix such a choice, and we omit $\varphi_{\mathbb{G}}$ from the notation, identifying $\mathbb{G}\simeq\mathbb{R}^{m}$. Note that $e\in\mathbb{G}$ is identified with $0\in\mathbb{R}^{m}$.
We recall from [Bel96, Prop. 5.18] the following result on the structure of the vector fields $X_{j}$.
Proposition 44.
In exponential coordinates $(x_{1},\dots,x_{m})\in\mathbb{R}^{m}$, the vector fields $X_{j}$, for $j=1,\dots,k$ have the following form:
(264) $X_{j}(x_{1},\dots,x_{m})=\partial_{j}+\sum_{h>k}Q_{jh}(x_{1},\dots,x_{k_{w_{h}-1}})\partial_{h},$
where $Q_{jh}:\mathbb{G}\simeq\mathbb{R}^{m}\to\mathbb{R}$ is a homogeneous polynomial of weighted degree $w_{h}-1$.
5.1. Endpoint maps
In this section we introduce the Endpoint map associated to a Carnot group. We work with the separable Hilbert space
(265) $H:=L^{2}(I,\mathbb{R}^{k}),$
for some fixed interval $I$, say $I=[0,1]$, and where $k$ is the rank of $\mathbb{G}$. For any $u\in H$ with $u=(u_{1},\dots,u_{k})$ we consider the following Cauchy problem for a curve $\gamma:I\to\mathbb{G}$:
(266) $\dot{\gamma}(t)=\sum_{j=1}^{k}u_{j}(t)X_{j}(\gamma(t)),\qquad\gamma(0)=e.$
It is well–known that (266) admits a unique absolutely continuous solution, $\gamma_{u}:I\to\mathbb{G}$, for any $u\in H$.
Definition 45.
Let $\mathbb{G}$ be a Carnot group. The Endpoint map of $\mathbb{G}$ is the function
(267) $\mathcal{E}:H\to\mathbb{G},$
given by $\mathcal{E}(u):=\gamma_{u}(1)$. We call the function $u\in H$ a control. We use the same notation to denote the Endpoint map $\mathcal{E}:H\to\mathbb{R}^{m}\simeq\mathbb{G}$ with an identification in exponential coordinates.
Thanks to the special form of the vector fields $X_{j}$ given by Proposition 44, we can rewrite the system (266) in exponential coordinates. Adopting the notation $\gamma=(\gamma_{1},\dots,\gamma_{m})$, we have
(268) $\begin{cases}\gamma_{i}(0)=0&i=1,\dots,m,\\ \dot{\gamma}_{i}(t)=u_{i}(t)&i=1,\dots,k,\\ \dot{\gamma}_{i}(t)=\displaystyle\sum_{j=1}^{k}u_{j}(t)Q_{ji}(\gamma_{1}(t),\dots,\gamma_{k_{w_{i}-1}}(t))&i=k+1,\dots,m,\end{cases}$
for a.e. $t\in I$. Similarly, we write $\gamma_{u}=(\gamma_{u,1},\dots,\gamma_{u,m})$ for the solution to (268).
The next result (Theorem H in the Introduction) connects this framework to the previous sections.
Theorem 46.
Let $\mathbb{G}$ be a Carnot group of topological dimension $m$, step $s$, and rank $k$. Then the Endpoint map $\mathcal{E}\in\mathscr{P}_{s}^{m}(H)$, for $H=L^{2}(I,\mathbb{R}^{k})$.
Proof.
The fact that $\mathcal{E}$ is $\mathcal{C}^{1}$ (in fact smooth) and that the differential $D\mathcal{E}$ is locally Lipschitz follows from [ABB20, Prop. 8.5]. The fact that the $D\mathcal{E}$ is weakly continuous is proved in [BL17, Thm. 23]. It remains to show that for all finite dimensional linear subspaces $E$ of $H$, the restriction $\mathcal{E}|_{E}:E\to\mathbb{R}^{m}$ is a polynomial map of degree at most $s$.
We actually prove a stronger claim: for any $n\in\mathbb{N}$, for any $n$-dimensional linear subspace $E\subset L^{2}(I,\mathbb{R}^{k})$, for any $t\in I$ and $i=1,\dots,m$, the function on $E$ defined by $E\ni u\mapsto\gamma_{u,i}(t)$ is a polynomial of degree $w_{i}$, and its coefficients (when $u$ is written in terms of some basis of $E\simeq\mathbb{R}^{n}$) are continuous functions of $t\in I$. We prove the claim by induction on $w$. We prove the base case, that is $w_{i}=1$. In this case $i=1,\dots,k$. For these values of $i$, by (268) we obtain
[…]
which is clearly a polynomial function of $u\in E$, of degree $1$. Moreover its coefficients (when $u$ is written in terms of a basis of $E\simeq\mathbb{R}^{n}$) are continuous functions of $t\in I$. This concludes the base case.
Let then $1\leq\theta<s$. The induction assumption is that for any $t\in I$ the function $E\ni u\to\gamma_{u,i}(t)$ is a polynomial of degree $w_{i}$ for all $i$ such that $w_{i}\leq\theta$, with coefficients that depends continuously on $t$.
Now we prove that the above property holds for any $i$ such that $w_{i}=\theta+1$. In particular $i>k$. In this case, by (268), the curve $\gamma_{u,i}:I\to\mathbb{R}$ is determined by
[…]
Recall that $Q_{ji}$ is a homogeneous polynomial of weighted degree $w_{i}-1$ and as already highlighted in the notation, depends only on the functions $\gamma_{u,\ell}(s)$ with $w_{\ell}\leq w_{i}-1=\theta$. By the induction assumption, all the functions $E\ni u\mapsto\gamma_{u,\ell}(s)$ are polynomials of degree $w_{\ell}\leq\theta$, depending continuously on $s$. By using the definition of weighted degree (see 4), it follows that the function
(271) $E\ni u\mapsto Q_{ji}(\gamma_{u,1}(s),\dots,\gamma_{u,k_{w_{i}-1}}(s)),$
is a polynomial of degree $w_{i}-1$, with coefficients depending continuously on $s\in I$. Then, the integrand in (270), namely the function
(272) $E\ni u\mapsto\sum_{j=1}^{k}u_{j}(s)Q_{ji}(\gamma_{u,1}(s),\dots,\gamma_{u,k_{w_{i}-1}}(s)),$
is a polynomial of degree $w_{i}=\theta+1$, with coefficients depending continuously on $s\in I$. We conclude easily using (270). ∎
Thanks to Theorem H we can apply the results of Section 4 to Endpoint maps of Carnot groups, obtaining Theorem I in the Introduction.
5.2. Sard property for real–analytic controls
We specialize the result in Theorem I to some concrete set of controls $K$, estimating its $n$–width.
Fix a closed interval $I\subset\mathbb{R}$ and let $\mathcal{D}(r,I):=\{z\in\mathbb{C}\mid d(z,I)\leq r\}$. The radius of convergence of a real–analytic function $u:I\to\mathbb{R}$ is the largest $r\in(0,\infty]$ such that $u$ extends to a (unique) holomorphic function on $\mathcal{D}(r,I)$. For fixed $r>0$ we denote by $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r)\subset L^{2}(I,\mathbb{R}^{k})$ the set of real–analytic controls whose components have radius of convergence strictly greater than $r$:
(273) $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r):=\bigg{\{}u\in L^{2}(I,\mathbb{R}^{k})\,\bigg{|}\,\textrm{$u$ is real--analytic with radius of convergence strictly greater than }r\bigg{\}},$
which we endow with the supremum norm:
(274) $\|u\|_{\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r)}:=\sup_{z\in\mathcal{D}(r,I)}\|u(z)\|.$
Similarly, we consider finite concatenations of real–analytic controls as above. More precisely, for $\ell\in\mathbb{N}$ divide the interval $I$ in equispaced sub-intervals:
(275) $I=I_{1}\cup\dots\cup I_{\ell},\qquad I_{a}=\inf I+\left[\frac{(a-1)|I|}{\ell},\frac{a|I|}{\ell}\right],$
and we let $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)\subset L^{2}(I,\mathbb{R}^{k})$ be the set of piecewise analytic controls $u:I\to\mathbb{R}^{k}$ that are real–analytic on each set $I_{1},\dots,I_{\ell}$, with radius of convergence strictly greater than $r$, namely
(276) $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)=\bigg{\{}u\in L^{2}(I,\mathbb{R}^{k})\,\bigg{|}\,u|_{I_{a}}\in\mathcal{C}^{\omega}(I_{a},\mathbb{R}^{k};r)\subset L^{2}(I_{a},\mathbb{R}^{k}),\,\text{ for all $a=1,\dots,\ell$}\bigg{\}},$
endowed with the norm
(277) $\|u\|_{\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)}=\max_{a=1,\dots,\ell}\|u\|_{\mathcal{C}^{\omega}(I_{a},\mathbb{R}^{k};r,\ell)}.$
Theorem 47.
Fix $I=[0,1]$, $\ell\in\mathbb{N}$ and $r>1$. Let $V=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$. Then the set $K_{V}:=\{u\in V\mid\|u\|_{V}\leq 1\}$ is compact in $L^{2}(I,\mathbb{R}^{k})$, and for its $n$–width it holds:
(278) $\Omega_{n}(K_{V},L^{2}(I,\mathbb{R}^{k}))\leq\frac{(k\ell)^{1/2}}{\ln r}\left(\frac{1}{r}\right)^{\left\lfloor\frac{n}{k\ell}\right\rfloor}.$
Proof.
We start by proving the case $\ell=1$. Fix $u=(u_{1},\dots,u_{k})\in K_{V}$. Since $r>1$, any $u_{j}$ for $j=1,\dots,k$ has an expansion in power series on the complex unit ball centered at the origin, denoted by $B_{\mathbb{C}}(1)$. In particular, we have for any $t\in I$
(279) $u_{j}(t)=\sum_{h=0}^{\infty}\frac{u^{(h)}_{j}(0)}{h!}t^{h}.$
For any $N\geq 1$ we consider the truncated sum, given for any $t\in I$ as
(280) $u_{j,N}(t)=\sum_{h=0}^{N-1}\frac{u^{(h)}_{j}(0)}{h!}t^{h}.$
By the Cauchy integral formula, and since $r>1$, we have
[…]
where $B_{\mathbb{C}}(r)$ is the complex unit ball of radius one centred at zero. Therefore,
(282) $|u_{j}^{(h)}(0)|\leq\frac{h!}{r^{h}}.$
Hence we have
[…]
It follows that for all $u\in K_{V}$ it holds
(284) $\|u-u_{N}\|_{L^{2}(I,\mathbb{R}^{k})}\leq\left(\sum_{j=1}^{k}\|u_{j}-u_{j,N}\|^{2}_{L^{2}(I)}\right)^{1/2}\leq\frac{k^{1/2}r^{-N}}{\ln r}.$
Thanks to this result we can find an approximating subspace. Consider the standard basis $e_{1},\dots,e_{k}$ of $\mathbb{R}^{k}$, and for any $k\in\mathbb{N}$ we denote by $f_{h}\in L^{2}(I)$ the function $f_{h}(t)=t^{h}$. We have
(285) $u_{N}=\sum_{h=0}^{N-1}\frac{u^{(h)}(0)}{h!}f_{h}=\sum_{j=1}^{k}\sum_{h=0}^{N-1}\frac{u^{(h)}_{j}(0)}{h!}f_{h}\otimes e_{j}.$
In particular, $u_{N}\in V_{N}:=\mathrm{span}\{f_{h}\otimes e_{j}\mid j=1,\dots,k,\;h=0,\dots,N-1\}$ and $\dim(V_{N})=kN$. Hence from (284) it follows that for any $N\in\mathbb{N}$
(286) $\Omega_{kN}(K_{V},L^{2}(I,\mathbb{R}^{k}))\leq\frac{k^{1/2}}{\ln r}\left(\frac{1}{r}\right)^{N}.$
For $n\in\mathbb{N}$ we consider $N=\left\lfloor\frac{n}{k}\right\rfloor$. By Definition 26, the $n$–width is non-increasing as a function of $n\in\mathbb{N}$, hence by (286) we have
(287) $\Omega_{n}(K_{V},L^{2}(I,\mathbb{R}^{k}))\leq\Omega_{kN}(K_{V},L^{2}(I,\mathbb{R}^{k}))\leq\frac{k^{1/2}}{\ln r}\left(\frac{1}{r}\right)^{\left\lfloor\frac{n}{k}\right\rfloor}.$
This proves the estimate on the $n$–width per $\ell=1$.
We sketch the argument for general $\ell$. In this case, for any $u\in\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$ we can write
(288) $u=\sum_{a=1}^{\ell}u|_{I_{a}}\mathbbm{1}_{I_{a}},$
where $I_{a}$ are the intervals of the decomposition (275), and $\mathbbm{1}_{I_{a}}$ are the corresponding characteristic functions. By definition, each $u|_{I_{a}}$ is the restriction to $I_{a}$ of a real–analytic function with radius of convergence $r_{a}\geq r>1$. We can expand each $u|_{I_{a}}$ in Taylor series centered at the lower bound of $I_{a}$. Then, in order to obtain the finite-dimensional approximation, we truncate the series as in the previous case, repeating analogous estimates for the remainder (taking into account the length of the intervals $|I_{a}|=1/\ell$). This concludes the estimate on the $n$–width for general $\ell$.
To prove the compactness assertion, observe that there is a linear immersion (given by the inclusion) $\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)\hookrightarrow L^{2}(I,\mathbb{R}^{k})$ and it holds $\|u\|_{L^{2}(I,\mathbb{R}^{k})}\leq\|u\|_{C^{\omega}(I,\mathbb{R}^{k};r,\ell)}$ for all $u\in C^{\omega}(I,\mathbb{R}^{k};r,\ell)$. In particular $K_{V}$ is bounded in $L^{2}(I,\mathbb{R}^{k})$ and since its $n$–width tends to zero as $n\to\infty$ it is also compact by Proposition 27. ∎
We can now prove Theorem J in the Introduction, of which we recall the statement.
Theorem 48.
Let $\mathbb{G}$ be a Carnot group of topological dimension $m$, step $s$, and rank $k$. Given $\ell\in\mathbb{N}$, there exists $r=r(m,s,k,\ell)>0$ such that, letting $V=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$, with $I=[0,1]$, it holds
(289) $\mu\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap V)\big{)}=\mu\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{V})\big{)}=0.$
Namely, the Sard property holds on the space of piecewise real–analytic controls with radius of convergence $>r$ and with $\ell$ pieces. Furthermore, letting $W=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};\infty,\ell)$, then it holds
(290) $\dim_{\mathcal{H}}\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E})\cap W)\big{)}\leq\dim_{\mathcal{H}}\big{(}\mathcal{E}(\mathrm{Crit}(\mathcal{E}|_{W})\big{)}\leq m-1.$
Namely, the strong Sard property holds on the space of piecewise entire controls with $\ell$ pieces.
Proof.
Observe that $\mathcal{E}\in\mathscr{P}^{m}_{s}(L^{2}(I,\mathbb{R}^{k}))$ by Theorem H. Let $V=\mathcal{C}^{\omega}(I,\mathbb{R}^{k};r,\ell)$, and $K_{V}$ be its unit ball with respect to the norm (277). By Theorem 47, for any $r>1$ we have
(291) $\limsup_{n\to\infty}\Omega_{n}(K_{V},L^{2}(I,\mathbb{R}^{k}))^{1/n}\leq r^{-\tfrac{1}{k\ell}}.$
We can apply Theorem I, observing that $\mathrm{span}(K_{V})=V$. ∎
5.3. Surjectivity of the Endpoint map on finite-dimensional spaces of controls
The Endpoint maps of sub-Riemannian manifolds are surjective when restricted to piecewise constant controls: this follows from the proof of the Rashevskii-Chow theorem in [ABB20]. In this section, we prove that in Carnot groups the Endpoint maps are surjective also when they are restricted to the space of controls which consists in the set of polynomial maps of some large enough fixed degree (which depends on the Carnot group). The proof is obtained by using a quantitative version of the inverse function theorem, as it can be found in [Cla76]. For completeness we present and prove in our setting the statements we need.
Given $M\in\mathcal{L}(\mathbb{R}^{m},\mathbb{R}^{m})$ we denote by $\sigma(M)$ the smallest singular value of $M$, namely
(292) $\sigma(M):=\inf_{\left\lVert v\right\rVert=1}\left\lVert Mv\right\rVert.$
From the definition it follows that for all $M_{1},M_{2}\in\mathcal{L}(\mathbb{R}^{m},\mathbb{R}^{m})$ it holds
(293) $|\sigma(M_{1})-\sigma(M_{2})|\leq\left\lVert M_{1}-M_{2}\right\rVert_{\operatorname{op}}.$
From [Cla76, Lemma 3] we obtain the following technical lemma.
Lemma 49.
Let $M_{0}\in\mathcal{L}(\mathbb{R}^{m},\mathbb{R}^{m})$ be invertible. Then, for every $v\in\mathbb{R}^{m}$ with $\|v\|=1$ there exists $w\in\mathbb{R}^{m}$ with $\|w\|=1$ such that
(294) $\langle w,Mv\rangle\geq\frac{\sigma(M_{0})}{2},$
for all $M$ such that $\|M-M_{0}\|_{\operatorname{op}}\leq\tfrac{\sigma(M_{0})}{2}$.
Proof.
Given $v\in\mathbb{R}^{m}$ with $\|v\|=1$, the set $C_{v}\subset\mathbb{R}^{m}$
(295) $C_{v}:=\left\{Mv\in\mathbb{R}^{m}\,\middle|\,\|M-M_{0}\|_{\operatorname{op}}\leq\frac{\sigma(M_{0})}{2}\right\},$
is convex. We now prove that its distance from 00 is at least $\frac{\sigma(M_{0})}{2}$. Indeed, by (293), for any $M$ with $\|M-M_{0}\|_{\operatorname{op}}\leq\tfrac{\sigma(M_{0})}{2}$ we obtain $\sigma(M)>\frac{\sigma(M_{0})}{2}$. Therefore,
(296) $\|Mv\|\geq\inf_{\|v\|=1}\|Mv\|=\sigma(M)>\frac{\sigma(M_{0})}{2}.$
Hence, we have proved that $C_{v}$ is separated from the ball $B_{\mathbb{R}^{m}}\left(\frac{\sigma(M_{0})}{2}\right)$. We conclude the proof thanks to the usual separation theorem for convex sets. Indeed, it directly provides $w\in\mathbb{R}^{m}$ with $\|w\|=1$ such that for all $M$ with $\|M-M_{0}\|_{\operatorname{op}}\leq\tfrac{\sigma(M_{0})}{2}$ we have
(297) $\langle w,Mv\rangle\geq\frac{\sigma(M_{0})}{2},$
concluding the proof. ∎
From [Cla76, Lemma 4] we obtain the following “quantitative injectivity” lemma.
Lemma 50.
Let $W\subset\mathbb{R}^{m}$ be an open set, and $f\in\mathcal{C}^{1}(W,\mathbb{R}^{m})$. Assume $x_{0}\in W$ is such that $D_{x_{0}}f$ is invertible. Set $\sigma_{f,x_{0}}:=\sigma(D_{x_{0}}f)>0$. There exists $r_{f,x_{0}}>0$ such that for any $g\in C^{1}(W,\mathbb{R}^{m})$ with
(298) $\sup_{z\in W}\left\lVert D_{z}f-D_{z}g\right\rVert_{\operatorname{op}}<\frac{\sigma_{f,x_{0}}}{4},$
and all $x,y\in B_{\mathbb{R}^{m}}(x_{0},r_{f,x_{0}})$ it holds
(299) $\|g(x)-g(y)\|\geq\frac{\sigma_{f,x_{0}}}{2}\|x-y\|.$
Proof.
Since $f$ is $\mathcal{C}^{1}$ there exists $r=r_{f,x_{0}}>0$ such that $B_{\mathbb{R}^{m}}(x_{0},r)\subset W$ and
(300) $\sup_{z\in B_{\mathbb{R}^{m}}(x_{0},r)}\left\lVert D_{z}f-D_{x_{0}}f\right\rVert_{\operatorname{op}}<\frac{\sigma_{f,x_{0}}}{4}.$
Let $g\in\mathcal{C}^{1}(W,\mathbb{R}^{m})$ as in the statement. By (298) we get
(301) $\left\lVert D_{z}g-D_{x_{0}}f\right\rVert_{\operatorname{op}}<\frac{\sigma_{f,x_{0}}}{2},$
for all $z\in B_{\mathbb{R}^{m}}(x_{0},r)$. Now we fix $x\neq y$ in $B_{\mathbb{R}^{m}}(x_{0},r)$, and we have
[…]
We define $v_{x,y}:=\frac{x-y}{\|x-y\|}\in\mathbb{R}^{m}$, and we apply Lemma 49 to $v=v_{x,y}$ and $M_{0}=D_{x_{0}}f$, which is invertible by hypothesis. We denote by $w_{x,y}\in\mathbb{R}^{m}$ the provided unit vector, such that
(303) $\left\langle w_{x,y},D_{z}g\cdot\frac{x-y}{\|x-y\|}\right\rangle\geq\frac{\sigma_{f,x_{0}}}{2},$
for all $z\in B_{\mathbb{R}^{m}}(x_{0},r)$. We conclude the proof by the Cauchy-Schwarz inequality, which gives
[…]
concluding the proof. ∎
We conclude with the following inclusion, analogous to [Cla76, Lemma 5], providing the “quantitative surjectivity” counterpart of the previous statement. We report the proof for completeness.
Lemma 51.
In the same setting of Lemma 50, the following inclusion holds
(305) $g(B_{\mathbb{R}^{m}}(x_{0},r_{f,x_{0}}))\supset B_{\mathbb{R}^{m}}\left(g(x_{0}),\frac{r_{f,x_{0}}\sigma_{f,x_{0}}}{8}\right).$
Proof.
Set $r=r_{f,x_{0}}$ and $\sigma=\sigma_{f,x_{0}}$. Let $y\in B_{\mathbb{R}^{m}}\left(g(x_{0}),\frac{r\sigma}{8}\right)$. Let $x$ be a minimum of $\|y-g(\cdot)\|^{2}$ on $B_{\mathbb{R}^{n}}(x_{0},r)$. We claim that $x\in\mathrm{int}B_{\mathbb{R}^{n}}(x_{0},r)$. Otherwise, using Lemma 50, we have
(306) $\frac{r\sigma}{8}$ $\geq\|y-g(x_{0})\|$
(307) $\geq\|g(x)-g(x_{0})\|-\|y-g(x)\|$
(308) $\geq\frac{\sigma}{2}\|x-x_{0}\|-\|y-g(x)\|$
(309) $\geq\frac{\sigma r}{2}-\|y-g(x_{0})\|$
(310) $\geq\frac{\sigma r}{2}-\frac{\sigma r}{8}=\frac{3r\sigma}{8},$
which is a contradiction. Thus $x$ yields a local minimum for the function $\|y-g(\cdot)\|^{2}$, and consequently $\nabla_{x}\|y-g(\cdot)\|^{2}=-2D_{x}g\cdot(y-g(x))=0$. Note that by (301) and the fact that the singular values are $1$-Lipschitz (see Lemma 17) it follows that $\sigma(D_{x}g)>\tfrac{\sigma}{2}$, hence $D_{x}g$ is invertible. Hence $y=g(x)$. ∎
We can now state and prove the surjectivity property of the Endpoint maps of Carnot groups.
Theorem 52.
For every Carnot group $\mathbb{G}$ there exists $d_{\mathbb{G}}\in\mathbb{N}$ such that
(311) ($\mathcal{E}\left(\{u\in L^{2}(I,\mathbb{R}^{k})\mid u\text{ is a polynomial map with components of degrees at most }d_{\mathbb{G}}\}\right)=\mathbb{G}.$
Proof.
The classical proof of the Rashevskii-Chow theorem provides $u_{0}\in L^{2}(I,\mathbb{R}^{k})$ such that $D_{u_{0}}\mathcal{E}$ is surjective and $\mathcal{E}(u_{0})=0$ (see e.g. [ABB20, Sec. 3.2]). In particular there exist $w_{1},\dots,w_{m}\in\ker(D_{u_{0}}\mathcal{E})^{\perp}$ such that the vectors $D_{u_{0}}\mathcal{E}(w_{1}),\dots,D_{u_{0}}\mathcal{E}(w_{m})$ are linearly independent. Now we consider the map $f:\mathbb{R}^{m}\to\mathbb{G}\simeq\mathbb{R}^{m}$ defined as
(312) $f(s_{1},\dots,s_{m})=\mathcal{E}\left(u_{0}+\sum_{i=1}^{m}s_{i}w_{i}\right).$
The differential $D_{0}f$ is invertible by construction, in particular $f$ covers a neighbourhood of $f(0)=0$. The idea of the proof is to construct a perturbation $g:\mathbb{R}^{m}\to\mathbb{G}\simeq\mathbb{R}^{m}$ of the form
(313) $g(s_{1},\dots,s_{m})=\mathcal{E}\left(q_{0}+\sum_{i=1}^{m}s_{i}p_{i}\right),$
where $q_{0},p_{1},\dots,p_{m}\in L^{2}(I,\mathbb{R}^{k})$ are suitably chosen polynomial maps, in such a way that the image of $g$ stills contains a neighbourhood of 00. This will be done by applying Lemma 51, as we now explain.
Given $\varepsilon_{1},\eta_{1}>0$ by the density of polynomial maps in $L^{2}(I,\mathbb{R}^{k})$ and by the continuity of $\mathcal{E}$ and $D\mathcal{E}$, there exists a polynomial map $q_{0}\in L^{2}(I,\mathbb{R}^{k})$ such that
(314) $\left\lVert\mathcal{E}(q_{0})\right\rVert<\varepsilon_{1}\qquad\text{ and }\qquad\left\lVert D_{u_{0}}\mathcal{E}-D_{q_{0}}\mathcal{E}\right\rVert_{\operatorname{op}}<\eta_{1}.$
Now we prove that we can find polynomial maps $p_{1},\dots,p_{m}$ and a neighbourhood $W\subset\mathbb{R}^{m}$ of 00 such that, setting $\sigma_{f,0}=\sigma(D_{x_{0}}f)$, it holds
(315) $\sup_{W}\left\lVert D_{s}f-D_{s}g\right\rVert_{\operatorname{op}}\leq\frac{\sigma_{f,0}}{4}.$
By the density of polynomial maps, given $\eta_{2}>0$ we consider $p_{1},\dots,p_{m}$ such that for any $i=1,\dots,m$
(316) $\left\lVert w_{i}-p_{i}\right\rVert<\eta_{2}.$
We get the following estimate
(317) $\left\lVert D_{0}f-D_{0}g\right\rVert_{\operatorname{op}}$ $\leq\sum_{i=1}^{m}\|D_{u_{0}}\mathcal{E}(w_{i})-D_{q_{0}}\mathcal{E}(p_{i})\|$
(318) $\leq\sum_{i=1}^{m}\|D_{u_{0}}\mathcal{E}(w_{i})-D_{q_{0}}\mathcal{E}(w_{i})\|+\|D_{q_{0}}\mathcal{E}(w_{i})-D_{q_{0}}\mathcal{E}(p_{i})\|$
(319) $\leq\left\lVert D_{u_{0}}\mathcal{E}-D_{q_{0}}\mathcal{E}\right\rVert_{\operatorname{op}}\sum_{i=1}^{m}\left\lVert w_{i}\right\rVert+\left\lVert D_{q_{0}}\mathcal{E}\right\rVert_{\operatorname{op}}\sum_{i=1}^{m}\left\lVert w_{i}-p_{i}\right\rVert$
(320) $\leq\eta_{1}\sum_{i=1}^{m}\left\lVert w_{i}\right\rVert+\left\lVert D_{q_{0}}\mathcal{E}\right\rVert_{\operatorname{op}}m\eta_{2}.$
Given $\eta_{3}>0$, by the continuity of $Df$ and $Dg$ there exists a neighbourhood $W\subset\mathbb{R}^{m}$ of 00 such that
(321) $\sup_{s\in W}\left\lVert D_{s}f-D_{0}f\right\rVert_{\operatorname{op}}+\left\lVert D_{s}g-D_{0}g\right\rVert_{\operatorname{op}}<\eta_{3}.$
Hence, by the triangle inequality we get the following estimate
(322) $\sup_{s\in W}\left\lVert D_{s}f-D_{s}g\right\rVert_{\operatorname{op}}\leq\eta_{3}+\eta_{1}\sum_{i=1}^{m}\left\lVert w_{i}\right\rVert+\left\lVert D_{q_{0}}\mathcal{E}\right\rVert_{\operatorname{op}}m\eta_{2}.$
To obtain (315) it is enough to choose $\eta_{1},\eta_{2},\eta_{3}$ small enough. We have thus found $g$ in such a way $\|f(0)-g(0)\|<\varepsilon_{1}$ and $g$ satisfies the hypothesis of Lemma 51. By the latter, we find $r_{f,0}>0$ such that
(323) $g(B_{\mathbb{R}^{m}}(r_{f,0}))\supset B_{\mathbb{R}^{m}}\left(g(0),\frac{r_{f,0}\sigma_{f,0}}{8}\right).$
Since $\|g(0)\|<\varepsilon_{1}$ by construction, taking $\varepsilon_{1}=\frac{r_{f,0}\sigma_{f,0}}{16}$ we also get
(324) $g(B_{\mathbb{R}^{m}}(r_{f,0}))\supset B_{\mathbb{R}^{m}}\left(\frac{r_{f,0}\sigma_{f,0}}{16}\right).$
Let then $d_{\mathbb{G}}$ be the maximum degree of the polynomial maps $q_{0},\dots,p_{m}$. We have proved that
(325) $B_{\mathbb{R}^{m}}\left(\frac{r_{f,0}\sigma_{f,0}}{16}\right)\subset\mathcal{E}\left(\{u\in L^{2}(I,\mathbb{R}^{k})\mid u\text{ is a polynomial map of degree at most }d_{\mathbb{G}}\}\right).$
In other words, the Endpoint map is surjective on a small ball when restricted to polynomial controls of degree $\leq d_{\mathbb{G}}$. Now recall that, for Carnot groups, dilations have the following property:
(326) $\delta_{\lambda}(\mathcal{E}(u))=\mathcal{E}(\lambda u),\qquad\forall\,\lambda>0,\quad u\in L^{2}(I,\mathbb{R}^{k}).$
From (325) and (326) it follows that $\mathcal{E}$, when restricted to polynomial controls of degree $\leq d_{\mathbb{G}}$ is surjective on the whole $\mathbb{G}\simeq\mathbb{R}^{m}$. ∎
With the same proof we can obtain the following statement, corresponding to Theorem K in the Introduction, of which Theorem 52 is a special case.
Theorem 53.
Let $\mathbb{G}$ be a Carnot group with topological dimension $m$ and rank $k$. Let $S\subset L^{2}(I,\mathbb{R}^{k})$ be a dense set. Then there exist $u_{0},u_{1},\dots,u_{m}\in S$ such that
(327) $\mathcal{E}(\mathrm{span}\{u_{0},u_{1},\dots,u_{m}\})=\mathbb{G}.$